### v2

# Moduli stacks of Higgs bundles on stable curves

Abstract
In this article, we construct a flat degeneration of the derived moduli stack of Higgs bundles on smooth curves using the stack of expanded degenerations of Jun Li. We show an intrinsic relative log-symplectic form on the degeneration, and we compare it with the one constructed by the second author. We show that the Hitchin map of the degeneration we construct has complete fibers. Furthermore, we show that the Hitchin map is flat and that a suitable open subset of the smooth locus of the reduced nilpotent cone is Lagrangian. We also extend the construction of the moduli of Higgs bundles along with the relative log-symplectic form over the universal moduli stack of stable curves.
1 Introduction
It is well known that moduli stacks of vector bundles on nodal curves are not complete. Often, their completions involve adding torsion-free sheaves as points of the boundary. Because we want to use the mapping space techniques of shifted symplectic geometry, we adopt the alternative approach of completing the moduli by adding boundary points, which are bundles on a bubbling node. This approach goes back to the classical works of David Gieseker [12] and Jun Li [17]. We will implement the approach in our setting by utilizing Jun Li’s stack of expanded degenerations. This allows us to use only vector bundles instead of torsion-free coherent sheaves and apply techniques of shifted symplectic geometry [32], obtaining a relative symplectic structure on a complete moduli stack. We extend the Hitchin map to Higgs bundles on the bubbled curves parametrized by the stack of expanded degenerations and then use the completeness of this moduli stack to prove that our extended Hitchin map is complete.
The derived moduli stack of Higgs bundles on a smooth curve $C$ has virtual dimension
$-\chi_{\text{Dol}}(C,\mathcal{E}nd(E,\phi))=n^{2}\deg(c_{1}(T_{C}))=2n^{2}(g-1)$
where $g$ is the genus of $C$ while the derived moduli stack of vector bundles (or any other Lagrangian in the derived moduli stack of Higgs bundle) has virtual dimension
$-\chi(Y,\mathcal{E}nd(E))=\int_{C}\text{ch}(\mathcal{E}nd(E))\text{td}(C)=n^{2}(g-1).$
These moduli stacks are homotopically locally of finite presentation, and so exhibit the “hidden smoothness” envisioned in derived deformation theory. Given this it is natural to study other geometric properties of these moduli. This article investigates relative 00-shifted symplectic structures on the total space of a degeneration of $C$ to a nodal curve.
After writing this article, we realized there are a lot of relations with the articles [2], [3] and significant overlap with a forthcoming article [11] of Ron Donagi and Andres Fernandez Herrero. In their work, they give a construction of a good moduli space for the semistable locus and semistable reduction relative to the Hitchin fibration (so the moduli space is proper over the Hitchin base in families). This uses the “infinite dimensional GIT” picture developed with Dan Halpern-Leistner. They also showed the flatness of the Hitchin morphism and that things are syntonic (as classical stacks), including the symplectic leaves. They are pursuing the story with punctures, thinking about the log Poisson reduction picture for the relative log cotangent stack of framed Gieseker vector bundles.
To start with, we fix a family of projective curves $\mathcal{X}$ over the spectrum of a discrete valuation ring $S$ such that the generic fibre is smooth, the closed fibre is an irreducible nodal curve with a single node, and the total space $\mathcal{X}$ is smooth. We also fix a rank $n$ and degree $d$ for the vector bundles in our moduli problem. In Section §2, we recall the construction of the stack $\mathfrak{M}$ of bounded expanded degenerations (bounded by the integer $n$) of the family of curves $\mathcal{X}/S$ following [17] and [42]. One of the main results in this subsection is Lemma 3.9, which we use to prove the following proposition.
Proposition 3.11  The morphism $\mathfrak{M}\longrightarrow S$ of ordinary Artin stacks is a log-smooth map. Moreover, the relative log-cotangent complex $\mathbb{L}^{log}_{\mathfrak{M}/S}=0$.
In subsection §2.3, we discuss shifted log-symplectic structures on quasi-smooth derived Artin stacks (Definition 3.4.0(a)) equipped with a locally free log structure. In subsection §2.4, we recall the definition of relative shifted symplectic forms for a quasi-smooth morphism of derived Artin stacks. We define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures.
In Section §3, we construct a logarithmic version of the relative Dolbeault moduli stack for the universal expanded family of curves $\mathcal{\mathcal{missing}}X_{\mathfrak{M}}\rightarrow\mathfrak{M}$. We show that the relative logarithmic Dolbeault moduli stack has a relative 00-shifted log-symplectic form over $S$. Moreover, we show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over $S$. This was proved for moduli schemes in [9]. The main results of this section are the following.
Theorem 4.7  $\mathcal{X}_{Dol}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $S$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol},BGL_{n}\times\mathcal{S})$ has a 00-shifted relative symplectic structure over $S$.
Theorem 4.9  The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathcal{X}_{\mathfrak{M},Dol},BGL_{n}\times\mathfrak{M})$ has a natural relative 00-shifted log-symplectic structure over $S$.
In Section §4, we define the Hitchin map on the classical Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see 8.1.2). We prove that the Hitchin map is complete.
Theorem 5.3  The morphism $h:M^{cl}_{Gie}\longrightarrow B$ is complete.
This result was proved in [1] for the Hitchin map on the moduli scheme where rank and degree are co-prime. We prove it here for the moduli stack, and the argument does not require us to assume that the rank and degree are coprime.
In section §5, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$, which is the reduction of the scheme theoretic fibre over the point $0\in B$. We prove that every irreducible component of of the reduced nilpotent cone has an open subset (denoted by $\mathcal{N}ilp^{sm,gen}$), which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat. The main theorem in this section is the following.
Theorem 6.9
  1. 1.
    The Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ is surjective.
  2. 2.
    The sub-stack $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic over $S$.
  3. 3.
    The Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ is flat.
In Section §6, we construct the Gieseker-like derived moduli stack of Higgs bundles $\mathcal{M}^{Dol}_{g}$ over the moduli stack of stable curves of genus $g\geq 2$. We prove the following theorem.
Theorem 7.6  There is a 00-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
In the Appendix, we construct the relative classical Artin stack of Gieseker-Higgs bundles and study its local properties. The main results of the appendix are Proposition 8.10 and Theorem 8.11. In the first lemma, we prove that the stack of Gieseker vector bundles is an almost very good stack. We use this in Lemma 8.11 to show that the classical stack of Gieseker-Higgs bundles is an irreducible, local complete intersection.
Proposition 8.10  The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([37, Definition 2.1.2]) with normal crossing singularities.
Theorem 8.11  The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $12\dim N_{Gie,0}+$ 1.
1.1 Notations and Conventions
  1. $\bullet$
    $\mathbb{k}$ be an algebraically closed field of characteristic zero. Every geometric object lives over $\operatorname{Spec}(\mathbb{k})$. However, the $\mathbb{k}$ is usually suppressed from the notation.
  2. $\bullet$
    $S:=\{\eta,o\}$ denotes the spectrum of a complete discrete valuation ring over $\mathbb{k}$, where $\eta$ denotes the generic point, and $o=\operatorname{Spec}(\mathbb{k})$ denotes the closed point.
  3. $\bullet$
    $\mathcal{X}\to S$ denotes a flat family of curves whose generic fibre is smooth projective, and the closed fibre is a nodal curve with a single node. We denote the nodal curve by $X_{0}$ and the node by $x$. We denote its normalisation by $q:\widetilde{X}_{0}\longrightarrow X_{0}$ and the two pre-images of the node $x$ by $\{x^{+},x^{-}\}$.
  4. $\bullet$
    $\mathbf{dg}$ is the category of dg-modules over $\mathbb{k}$ (i.e. of complexes of $\mathbb{k}$-modules). By convention, the differential of an object in $\mathbf{dg}$ increases degrees. So an object is a cochain complex of the form
    $\cdots\to E_{-1}\to E_{0}\to E_{1}\to\cdots.$
  5. $\bullet$
    $\mathrm{cdga}$ is the category of commutative dg-algebras over $\mathbb{k}$, and $\mathrm{cdga}^{\leq 0}$ its full subcategory of non-positively graded commutative dg-algebras.
  6. $\bullet$
    $\mathbf{dg}$, $\mathrm{cdga}$ (respectively $\mathrm{cdga}^{\leq 0}$) are endowed with their natural model structures for which equivalences are quasi-isomorphisms, and fibrations are epimorphisms (respectively, epimorphisms in strictly negative degrees).
  7. $\bullet$
    ${\operatorname{dAff}}:=(\mathrm{cdga}^{\leq 0})^{op}$ is the category of derived affine $\mathbb{k}$-schemes.
  8. $\bullet$
    The $\infty$-categories associated with the model categories $\mathbf{dg},\mathrm{cdga}^{\leq 0},\operatorname{dAff}$ are denoted by $\textbf{dg},\textbf{cdga}^{\leq 0},\textbf{dAff}$.
  9. $\bullet$
    The $\infty$-category of simplicial sets is denoted by $\mathbb{S}$. It is also called the $\infty$-category of spaces, and space will be used to mean a simplicial set.
  10. $\bullet$
    The $\infty$-category of derived stacks over $\mathbb{k}$, for the étale topology, is denoted by dSt. If $X$ is a derived stack, the $\infty$-category of derived stacks over $X$ is denoted by $\textbf{dSt}_{X}$. The truncation of a derived stack $X$ is denoted by $t_{0}X$.
  11. $\bullet$
    Our running convention is that unless explicitly specified otherwise a (derived) Artin stack will always mean a (derived) geometric $1$-stack. In particular, all the moduli stacks we consider are geometric $1$-stacks.
  12. $\bullet$
    List notations for log structures and classifying stacks of log structures.
  13. $\bullet$
    For a family of semistable curves $\mathfrak{X}$ over a scheme $T$, $\mathfrak{X}_{Dol}$ denotes the relative logarithmic Dolbeault shape of $\mathfrak{X}/T$.
    ${\mathfrak{X}_{Dol}}$${\mathfrak{X}}$${T}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$
2 Acknowledgements
T.P. (University of Pennsylvania) was partially supported by NSF FRG grant DMS- 2244978, NSF/BSF grant DMS-2200914, NSF grant DMS-1901876, and Simons Collaboration grant number 347070. O.B. (University of Haifa) and S.D. would like to acknowledge the Binational Science Foundation and NSF/BSF grant 2021717 for supporting S.D. as a postdoc. S.D. would like to thank Professor Alek Vainshtein of the University of Haifa, Israel, for the financial support from his Israel Science Foundation grant number 876/20 during a postdoc position at the University of Haifa. S.D. would like to thank Professor Vikraman Balaji for the financial support from his SERB Core Research Grant CRG/2022/000051-G during a postdoc position at Chennai Mathematical Institute. S.D. would also like to thank IISER Tirupati for their support during a part of this work.
3 Preliminaries
3.1 Space of bounded Expanded degenerations
In this subsection, we will recall a construction by Jun Li [17] called the stack of expanded degenerations. We will use this stack to construct our degeneration. The construction of Jun Li starts with a one-parameter family of varieties (of any dimension) such that the total space of the family is smooth, the generic fibre is smooth, and the special fibre is a normal crossing divisor in the total space of the family. For our purpose, we only need to consider the expanded degenerations of such a family for curves. We will make use of the standard fact that for any nodal curve, one can always construct a smoothing over the spectrum of a discrete valuation ring, which also has a smooth total space. The precise setting we consider can be spelled out as follows.
3.1.1 Degeneration of curves.
Start with a flat family of projective curves $\mathcal{X}\longrightarrow S$, such that
  1. 1.
    the generic fibre $\mathcal{X}_{\eta}$ is a smooth curve of genus $g\geq 2$,
  2. 2.
    the closed fibre is a nodal curve $X_{0}$ with a single node $x\in X_{0}$, and
  3. 3.
    the total space $\mathcal{X}$ is regular over ${\rm Spec}\,\mathbb{k}$.
Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. From here onwards, we assume that $S$ is a neighbourhood of the origin in $\mathbb{A}^{1}:=Spec\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{k}[t]$.
Definition 3.1.
(Gieseker curve/Expanded degeneration/Modification) Let $X_{0}$ be a nodal curve with a single node $x\in X_{0}$, and let $x^{\pm}$ label the two preimages of $x$ in the normalization $\widetilde{X}_{0}$ of $X_{0}$. Let $r$ be a positive integer.
  1. 1.
    A chain of $r$ projective lines is a scheme $R[r]$ of the form $\cup_{i=1}^{{}^{r}}R[r]_{i}$ such that
    1. (a)
      $R[r]_{i}\cong\mathbb{P}^{{}^{1}}$,
    2. (b)
      for any $i<j$, $R[r]_{i}\cap R[r]_{j}$ consists of a single point $p_{j}$ if $j=i+1$ and empty otherwise.
    We call $r$ the length of the chain $R[r]$. Let us choose and fix two smooth points $p_{1}$ and $p_{r+1}$ on $R[r]_{1}$ and $R[r]_{r}$, respectively.
    $p_{1}$$p_{r+1}$
  2. 2.
    A Gieseker curve $X_{r}$ is the categorical quotient of the disjoint union of the curves $\tilde{X}_{0}$ and $R[r]$ obtained by identifying $x^{+}$ with $p_{1}$ and $x^{-}$ with $p_{r+1}$. We will also refer to these curves as expanded degenerations or bubblings of the nodal curve $X_{0}$.
Definition 3.2.
(Family of Gieseker curves/Expanded degenerations/Modifications:) For every $S$-scheme $T$, a modification of $\mathcal{X}$ relative to $T$ is a commutative diagram
${\mathfrak{X}_{T}}$${\mathcal{X}_{T}:=\mathcal{X}\times_{S}T}$${T}$$\scriptstyle{p_{T}}$$\scriptstyle{\pi_{T}}$ (3.1)
such that
  1. 1.
    $p_{T}:\mathfrak{X}_{T}\longrightarrow T$ is flat;
  2. 2.
    the horizontal morphism is finitely presented and is an isomorphism over smooth fibers $(\mathcal{X}_{T})_{t}$ of $\mathcal{X}_{T}\to T$;
  3. 3.
    over each closed point $t\in T$ which maps to $\eta_{0}\in S$, we have $(\mathfrak{X}_{T})_{t}\cong X_{r}$ for some integer $r$ and the horizontal morphism restricts to the morphism $X_{r}\to X_{0}$ which contracts the $\mathbb{P}^{1}$’s in $X_{r}$ to the node $x\in X_{0}$.
Definition 3.3.
(Morphisms of families of expanded degenerations:) Let $T\to S$ be a scheme mapping to $S$, and let $\mathfrak{X}_{T}$ and ${\mathfrak{X}^{\prime}}_{T}$ be two $T$-relative modifications of $\mathcal{X}\to S$.
We call $\mathfrak{X}_{T}$ and $\mathfrak{X}^{\prime}_{T}$ isomorphic if there exists an isomorphism $\sigma_{T}:\mathfrak{X}_{T}\longrightarrow\mathfrak{X}^{\prime}_{T}$ such that the following diagram commutes
${\mathfrak{X}_{T}}$${\mathfrak{X}^{\prime}_{T}}$${\mathcal{X}_{T}}$$\scriptstyle{\pi_{T}}$$\scriptstyle{\sigma_{T}}$$\scriptstyle{\pi^{\prime}_{T}}$ (3.2)
3.1.2 Space of expanded degenerations
As before we will choose an uniformizer on $S$ so that $S$ is identified with a neighborhood of zero in $\mathbb{A}^{1}$. For any positive integer $n$, we set
$S[n]:=S\times_{\mathbb{A}^{1}}\mathbb{A}^{n+1}$ (3.3)
where the map $\mathbb{A}^{n+1}\longrightarrow\mathbb{A}^{1}$ is given by $(t_{1},\dots,t_{n+1})\mapsto t_{1}\cdots t_{n+1}$.
Consider the group
$G[n]:=\mathbb{G}_{m}^{n}$ (3.4)
acting on $\mathbb{A}^{n+1}$ by
$(\sigma_{1},\ldots,\sigma_{n})\cdot(t_{1},\ldots,t_{n+1}):=(\sigma_{1}\cdot t_{1},\ldots,\sigma_{i-1}^{-1}\cdot\sigma_{i}\cdot t_{i},\ldots,\sigma_{n}^{-1}\cdot t_{n+1}).$ (3.5)
The map $\mathbb{A}^{n+1}\to\mathbb{A}^{1}$, $(t_{1},\ldots,t_{n+1})\mapsto t_{1}t_{2}\cdots t_{n+1}$ intertwines this action with the trivial action on $\mathbb{A}^{1}$ and hence (3.5) induces an action of $G[n]$ on $S[n]$.
In [17, Section 1.1], Jun Li constructed a $G[n]$-equivariant family of expanded degenerations $W[n]$ over $S[n]$. It is constructed by several birational transformations of the family $\mathcal{X}\times_{S}S[n]$. The curves occurring in the family are all possible expanded degenerations of the family $\mathcal{X}/S$ for which the length of rational chains is bounded by $n$.
3.1.3 Stack of bounded Expanded degenerations
The significance of the family $W[n]\to S[n]$ is that it can be used as an atlas for the universal family of expanded degenerations with bubbling of length $\leq n$.
Indeed, recall from [17, Definition 1.9, Proposition 1.10] the following
Definition 3.4.
The stack of expanded degenerations of $\mathcal{X}/S$ is the stack
$\mathfrak{M}:{Sch}/S\longrightarrow{Groupoids}$
given by the assignment
[…]
Two such families $W_{T}$ and $W^{\prime}_{T}$ are isomorphic if there is a $T$-isomorphism $f:W_{T}\longrightarrow W^{\prime}_{S}$ compatible to the tautological projections $W_{T}\longrightarrow W$ and $W^{\prime}_{T}\longrightarrow W$.
By construction, the stack of expanded degenerations has the following properties
  1. 1.
    $\mathfrak{M}$ is a smooth Artin stack of finite type.
  2. 2.
    The projection map $\mathfrak{M}\longrightarrow S$ is generically an isomorphism.
  3. 3.
    The closed fibre of $\mathfrak{M}\to S$ is a normal crossing divisor in $\mathfrak{M}$.
Here we use the following
Definition 3.5.
Let $\mathcal{Y}$ be a smooth Artin stack and $\mathcal{D}$ be a closed sub-stack of co-dimension one. We say that $\mathcal{D}$ is a normal crossing divisor in $\mathcal{Y}$ if for any smooth morphism $f:T\rightarrow\mathcal{Y}$ from a smooth scheme $T$, the pull-back of the divisor $\mathcal{D}\times_{\mathcal{Y}}T$ is a normal crossing divisor in $T$.
3.1.4 An alternative construction of the stack $\mathfrak{M}$
In [42, Definition 2.22], Zijun Zhou gave an alternative construction of the stack $\mathfrak{M}$ which will be useful for our purpose. Here we briefly recall his construction.
First let us introduce some useful notation and terminology. For any integer $n$, we set $[n]:=\{1,\dots,n+1\}$. It will be useful to have short names for the natural maps between the spaces $S[k]$ for various values of $k$.
Definition 3.6.
Given any subset $I\subseteq[n]$ of cardinality $k+1$, we define
  • The Standard Embedding corresponding to $I$ is the embedding
    $\gamma_{I}:S[k]\hookrightarrow S[n]$ (3.7)
    given by $\gamma_{I}(t_{1},\dots,t_{k+1})=t_{i}$ if $i\in I$ and $\gamma_{I}(t_{1},\dots,t_{k+1})=1$ otherwise.
  • The Standard Open Embedding corresponding to $I$ is the open embedding
    $\tau_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\hookrightarrow S[n]$ (3.8)
    where $S[k]\times\mathbb{G}_{m}^{n-k}$ denotes the open subset swept by $S[k]$ under the free action of the subgroup $\mathbb{G}_{m}^{n-k}\subset G[n]$ whose $j$-th component is $1$ if $j\notin I$.
Definition 3.7.
Given two subsets $I$ and $I^{\prime}$ of $[n]$ both of order $k+1$, define an equivalence relation $R_{I,I^{\prime}}$ on $S[n]$ by setting
$R_{I,I^{\prime}}=S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$ (3.9)
where the two maps are $\tau_{I}$ and $\tau_{I^{\prime}}$.
For every $k\leq 0$ define a discrete equivalence relation $R_{k}\rightrightarrows S[n]$ as
$R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}R_{I,I^{\prime}}\ \rightrightarrows\ S[n].$ (3.10)
Define the amalgamated discrete equivalence relation $R_{dis}$ on $S[n]$ by setting
$R_{dis}:=\coprod_{k\leq n}R_{k}\ \rightrightarrows\ S[n]$ (3.11)
Finally we define the total smooth groupoid
$R_{tot}:=\mathbb{G}_{m}^{n}\times R_{dis}\rightrightarrows S[n]$
which is generated by $R_{dis}$ and the action of $\mathbb{G}_{m}^{n}$ on $S[n]$.
Definition 3.8.
Let $\mathfrak{M}_{n}:=\left[S[n]/R_{tot}\right]$ be the quotient stack of $S[n]$ by the smooth groupoid $R_{tot}\rightrightarrows S[n]$.
From [42, Remark 2.23], it follows that the stack $\mathfrak{M}_{n}$ is the stack of expanded degenerations of the family $\mathcal{X}/S$ bounded by the integer $n$.
Notation:  From here onwards, we will only work with $\mathfrak{M}_{n}$ and we drop the subscript "$n"$ from the notation and denote it simply by $\mathfrak{M}$.
Lemma 3.9.
The morphism between quotient stacks
$[S[n]/G[n]]\longrightarrow\mathfrak{M}$ (3.12)
is étale.
Proof.
Let us first show that the relation $R_{dis}$ descends to an equivalence relation on the quotient stack $[S[n]/G[n]]$). Recall that $R_{dis}:=\underset{k\leq n}{\coprod}R_{k}$ and
$R_{k}:=\underset{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}{\coprod}R_{I,I^{\prime}},$
where the equivalence relations $R_{I,I^{\prime}}$ are given by
$S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$ (3.13)
Recall that the map $R_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\longrightarrow S[n]$ is given by the standard open embedding.
For simplicity, let us assume that $S=\mathbb{A}^{1}$. Then $S[n]=\mathbb{A}^{n+1}$. Notice that $G[n]:=\mathbb{G}_{m}^{n}$. But in fact, there is an action of a bigger group $\mathbb{G}_{m}^{n+1}$ on $\mathbb{A}^{n+1}$ given by the coordinatewise multiplication. Moreover, the group $\mathbb{G}_{m}^{n}$ can be identified as a subgroup of $\mathbb{G}_{m}^{n+1}$ consisting of elements of the form $(t_{1},\dots,t_{n+1})$ such that $\prod^{n+1}_{i=1}t_{i}=1$. Let’s consider the action of the bigger group $G[n+1]$.
Let us first define a natural action of $G[n+1]$ on $R_{I}:=\mathbb{A}^{k+1}\times\mathbb{G}_{m}^{n-k}$. Notice that the cardinality of the set $I$ equals $k+1$. It is an ordered subset of $[n+1]$. Therefore, the complement $J:=[n+1]\setminus I$ is also an ordered set of cardinality $n-k$, where the order is the induced order from $[n+1]$. Now we see that $\mathbb{G}_{m}^{n+1}\cong\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. Finally, we define the action of $\mathbb{G}_{m}^{n+1}$ on $R_{I}:=\mathbb{A}^{I}\times\mathbb{G}_{m}^{J}$ by the obvious action of $\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. It is easy to check that the map $R_{I}:\mathbb{A}^{k+1}\times(\mathbb{k}^{*})^{n-k}\longrightarrow\mathbb{A}^{n+1}$ is $\mathbb{G}_{m}^{n+1}$-equivariant. Hence, it is also equivariant under the action of the smaller subgroup $\mathbb{G}_{m}^{n}$. Therefore, the equivalence relations $R_{I,I^{\prime}}$ descends to the quotient stack $[S[n]/G[n]]$.
Now we will show that the relation $\underline{R}_{dis}$ is an étale equivalence relation. First of all, notice that the maps $R_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\longrightarrow S[n]$ are open immersions. Therefore, $R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}R_{I,I^{\prime}}\rightrightarrows\mathbb{A}^{k+1}$ is an étale equivalence relation because both the projections are Zariski open immersions. Therefore, $R_{dis}$ on $S[n]$ and $[R_{dis}/G[n]]$ on $[S[n]/G[n]]$ defines an étale equivalence relation. ∎
3.2 Construction of the family of curves
In this subsection, we will briefly recall from [17, 42] the fact that there is a family of expanded degeneration (whose rational chain has length bounded by $n$). Let us choose an étale neighborhood $U_{p}$ of the node $p\in\mathcal{X}$ such that the nodal curve $U_{p}\times_{\mathcal{X}}X_{0}$ is a reducible nodal curve with two smooth connected components intersecting transversally at the node $p$. Let $V=\mathcal{X}-\{p\}$. Then $\{U_{p},V\}$ is an étale covering of $\mathcal{X}$. We have the following diagram (which is both cartesian and co-cartesian/push-out):
${U_{p}\times_{\mathcal{X}}V}$${V}$${U_{p}}$${\mathcal{X}}$ (3.14)
Notice that $U_{p}\longrightarrow S$ is a simple normal crossing degeneration which we can feed into the construction of [17, 42] to produce a family of expanded degenerations. We denote this family by $U_{p}[n]\longrightarrow S[n]$ and also write $V[n]:=V\times_{S}S[n]$. Now we construct the family of expanded degenerations, denoted by $\mathcal{X}[n]$ over $S[n]$ by the following push-out
${U_{p}[n]\times_{(\mathcal{X}\times_{S}S[n])}V[n]}$${V[n]}$${U_{p}[n]}$${\mathcal{X}[n]}$ (3.15)
The push-out exists as an algebraic space by Rydh’s pushout theorem [34, Theorem C]. The family of curves $\mathcal{X}[n]\longrightarrow S[n]$ is the desired family of expanded degenerations of the original family $\mathcal{X}\longrightarrow S$. We refer to [42, Proposition 2.13] for the key properties of the family of expanded degenerations. In particular, by properties (3) and (4) of [42, Proposition 2.13] the family $U_{p}[n]\to S[n]$ is equivariant under the action of the smooth groupoid $R_{tot}\rightrightarrows S[n]$. By the universality of pushouts this implies that the family of expanded degenerations $\mathcal{X}[n]\to S[n]$ is also equivariant and so $\mathcal{X}[n]$ descends to the stack of expanded degenerations $\mathfrak{M}=\mathfrak{M}_{n}=\left[S[n]/R_{tot}\right]$. We denote this family of expanded degenerations by $\mathfrak{X}_{\mathfrak{M}}\to\mathfrak{M}$.
3.3 Log structures on derived Artin stacks
In this subsection, we discuss shifted log-symplectic structure on a log quasi-smooth derived Artin stack (Definition 3.4.0(b)) equipped with a locally free log structure. We refer to [26, 27] and [29] for the prerequisite material on log-structure/locally free log-structure on Artin/derived Artin stacks. We now recall a few necessary definitions.
Let $\mathfrak{L}^{0}$ denote the algebraic stack which classifies fine log structures, and let $\mathfrak{L}^{1}$ denote the algebraic stack which classifies morphisms of fine log schemes [27]. We denote by $\mathfrak{L}^{0}_{\mathsf{lf}}$ and $\mathfrak{L}^{1}_{\mathsf{lf}}$ the classifying substacks (known to be smooth by [29, Proposition 1.7]) of locally free log structures and morphisms between locally free log structures, respectively. Given a logarithmic scheme $(S,P,\alpha)$, the stack $\mathfrak{L}^{0}_{S,\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{L}^{0}_{\mathsf{lf}}}S$ is the classifying stack of logarithmic morphisms from a scheme with locally free log structure to $(S,P,\alpha)$. With this notation, we now have the following natural definition [29]:
Definition 3.10.
A locally free log structure on a derived algebraic stack $\mathfrak{X}$ is a morphism $\mathfrak{X}\longrightarrow\mathfrak{L}^{0}_{\mathsf{lf}}$ of derived stacks. A morphism of derived log stacks $\mathfrak{X}\longrightarrow\mathfrak{Y}$ equipped with locally-free log structures is a commutative diagram of derived stacks
${\mathfrak{X}}$${\mathfrak{L}^{1}_{\mathsf{lf}}}$${\mathfrak{Y}}$${\mathfrak{L}^{0}_{\mathsf{lf}}}$ (3.16)
3.3.1 Relative logarithmic cotangent complex
Given a log morphism $f:\mathfrak{X}\longrightarrow\mathfrak{Y}$ between derived stacks equipped with locally free log structures, the relative log-cotangent complex of the morphism $f$ is defined by
${\mathbb{L}}^{log}_{\mathfrak{X}/\mathfrak{Y}}:=\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$ (3.17)
Here, as before, $\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{Y}}\mathfrak{L}^{0}_{\mathsf{lf}}$ and $\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$ denotes the relative cotangent complex for the morphism of stacks $\mathfrak{X}\longrightarrow\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}$.
With this definition in mind, we return to the discussion on the stack of bounded expanded degenerations $\mathfrak{M}\rightarrow S$ (3.4, 3.8) for a given one parameter degeneration $\mathcal{X}\rightarrow S$ of a smooth projective curve to a nodal curve.
Proposition 3.11.
The morphism $\mathfrak{M}\longrightarrow S$ of ordinary Artin stacks is a log-smooth map. Moreover, the relative log-cotangent complex $\mathbb{L}^{log}_{\mathfrak{M}/S}=0$.
Proof.
From Lemma 3.9, it follows that the map $[S[n]/G[n]]\longrightarrow\mathfrak{M}$ is étale. Notice that the varieties $S[n]$ and $S$ are log smooth varieties with the log structure coming from the divisor given by the pre-image of the closed point of $S$ under the map $S[n]\longrightarrow S$. Since the map $S[n]\longrightarrow S$ is an étale base-change of the map $\mathbb{A}^{n+1}\longrightarrow\mathbb{A}^{1}$ given by $(t_{1},\dots,t_{n+1})\mapsto t_{1}\cdots t_{n+1}$, the map is a log-smooth morphism. Therefore, it follows that the map $[S[n]/G[n]]\longrightarrow S$ is also a log-smooth morphism of log-smooth Artin stacks.
The explicit description of the relative log-cotangent complex is the following. The pull-back (to $S[n]$) of the log-cotangent complex of the stack $[S[n]/G[n]]$ is the perfect complex on $S[n]$ concentrated in degrees 00 and $1$ which is given by
$\left[\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$ (3.18)
where $\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])$ is the rank $(n+1)$ locally free sheaf of logarithmic one forms on $S[n]$ with poles along the strict normal crossings divisor $\partial S[n]=S[n]\times_{S}\{o\}$. Since the action of $G[n]$ on $S[n]$ respects $\partial S[n]$, we get a natural $G[n]$-equivariant structure $\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])$ such that the natural log-symplectic structure is $G[n]$-equivariant. The morphism (3.18) is given by the (equivariant) moment map for the action of $G[n]$ on the log-cotangent bundle of $S[n]$. In particular (3.18) is a $G[n]$-perfect complex. We recall that the action of $G[n]$ on $S[n]$ is given by
$(t_{1},\dots,t_{n+1})\cdot(x_{1},\dots,x_{n+1})=(t_{1}x_{1},t_{2}x_{2},\dots,t_{n+1}x_{n+1})$ (3.19)
where $t_{1}\cdots t_{n+1}=1$, and that the map $S[n]\longrightarrow S$ is given by
$(t_{1},\dots,t_{n+1})\mapsto\prod^{n+1}_{i=1}t_{i}$ (3.20)
It is easy to see that the map 3.18
$\left[\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$ (3.21)
of vector bundles is surjective: consider the action of $(\mathbb{G}_{m})^{n}$ on $\mathbb{A}^{n+1}$ given by 3.19. Let $p:=(x_{1},\cdots,x_{n+1})\in\mathbb{A}^{n+1}$ be any point. Then consider the orbit map at the point $p$.
$Or_{p}:(\mathbb{G}_{m})^{n}\longrightarrow\mathbb{A}^{n+1}$ (3.22)
$(t_{1},\cdots,t_{n+1})\mapsto(t_{1}\cdot x_{1},\cdots,t_{n+1}\cdot x_{n+1})$ (3.23)
The induced map on the tangent spaces is $d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}\longrightarrow T_{p}\mathbb{A}^{n+1}=\mathbb{C}^{n+1}$ is given by the multiplication by the following diagonal matrix
$\begin{bmatrix}x_{1}&0&0&\dots&0\\ 0&x_{2}&0&\dots&0\\ {5}\\ 0&0&0&\dots&x_{n+1}\end{bmatrix}$
This implies that the image of $[\sum^{n+1}_{i=1}\lambda_{i}\partial_{t_{i}}]\in T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}$ (notice that $\sum^{n+1}_{i=1}\lambda_{i}=0$) is $\sum^{n+1}_{i=1}\lambda_{i}\cdot x_{i}\cdot\partial_{x_{i}}=\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{log\leavevmode\nobreak\ x_{i}}$. Therefore, we get a map
$d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}\longrightarrow(T\mathbb{A}^{n+1}(-log\leavevmode\nobreak\ \partial A^{n+1}))_{p}=\mathbb{C}^{n+1}$ (3.24)
$(\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{t_{i}})\mapsto\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{log\leavevmode\nobreak\ x_{i}}$ (3.25)
The above map is clearly injective. Therefore the dual of this map
$(d(Or)_{p})^{\vee}:\Omega^{1}_{p}\mathbb{A}^{n+1}(log\leavevmode\nobreak\ \partial A^{n+1})\longrightarrow\Omega^{1}_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}$ (3.26)
is surjective. This is the explicit description of the moment map 3.18. It is easy to see that the composite map
$\Omega_{S}^{1}(\log o)\hookrightarrow\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ (3.27)
is 00, and that the morphism $\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism.
Notice that the perfect complex
$\big{[}\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\big{]}$ (3.28)
is precisely the relative log cotangent complex of the morphism $\big{[}S[n]/G[n]\big{]}\longrightarrow S$ pulled back to $S[n]$. But since the morphism $\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism, the relative log-cotangent complex of the morphism $\big{[}S[n]/G[n]\big{]}\longrightarrow S$ is equivalent to 00. ∎
3.4 Relative shifted log-symplectic forms
In this subsection, we will recall the definition of relative shifted symplectic forms for a l.f.p. morphism of derived Artin stacks. We will then define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures.
Let $\pi:M\longrightarrow S$ be a morphism of derived Artin stacks with $S$ a scheme. Suppose we also assume that the stack $M$ and the scheme $S$ are equipped with locally free log structures such that the map $\pi$ is a morphism of log stacks. These data are equivalent to a map of derived stacks $\pi_{log}:M\longrightarrow\mathfrak{L}^{0}_{S,\mathsf{lf}}$, where $\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is the classifying log stack for locally free log schemes mapping into $S$. We further assume that the map $\pi_{log}$ is of locally of finite presentation.
Recall that clasically a map $R\longrightarrow S$ of discrete commutative $\mathbb{k}$-algebras is finitely presented if $\mathsf{Hom}_{R/calg}(S,-)$ commutes with filtered colimits. Above we used the homotopical notion of being locally finitely presented, defined as follows:
Definition 3.12.
[33, Definition 2.16]  We say that a map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is (homotopically) finitely presented if $\mathsf{Map}_{A/\mathrm{cdga}^{\leq 0}}(B,-)$ commutes with (homotopy) filtered colimits.
Remark 3.13.
A map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is finitely presented ([35, 2.2.] and [22, Theorem 7.4.3.18]) if and only if
  1. 1.
    $H^{0}(A)\longrightarrow H^{0}(B)$ is classically finitely presented, and
  2. 2.
    $\mathbb{L}_{B/A}$ is a perfect $B$-dg module (i.e., a dualizable object in $(dgmod(B),\otimes)$).
We will also need an appropriate notion of quasi-smoothness which we recall next:
Definition 3.14.
  1. (a)
    A map of derived Artin stacks $f:\mathcal{X}\longrightarrow\mathcal{Y}$ is quasi-smooth if it is locally of finite presentation and the relative cotangent complex $\mathbb{L}_{f}$ concentrated in degrees $\geq-1$.
  2. (b)
    Let $\pi:M\rightarrow S$ be a logarithmic morphism from a derived Artin stack $M$ to a scheme $S$ both equipped with locally free log structures. We say that the map $\pi$ is log quasi-smooth if the corresponding classifying morphism $\pi_{log}:M\to\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is quasi-smooth.
Definition 3.15.
[29, Definition 2.12] The relative log cotangent complex of the log morphism of derived stacks $\pi:M\to S$ is the complex ${\mathbb{L}}^{log}_{M/S}:=\mathbb{L}_{\pi_{log}}$. Here $\mathbb{L}_{\pi_{log}}$ denotes the relative cotangent complex for the classifying morphism $\pi_{log}$.
Remark 3.16.
When $\pi:M\to S$ is a log quasi-smooth morphism from a derived Artin stack to a scheme $X$, then the quasi-smoothness of the map $\pi_{log}:M\longrightarrow\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is equivalent to the fact that ${\mathbb{L}}^{log}_{M/S}$ is a perfect complex over $M$ of amplitude $[-1,1]$.
Now we recall the definitions of relative shifted symplectic forms for a finitely presented map of derived Artin stacks $M\longrightarrow N$ from [32], [8, Definition 1.4.1]. The finitely presented condition implies that the relative cotangent and relative tangent complexes are perfect (Remark 3.13).
Let $A\in\mathrm{cdga}^{\leq 0}$ and let $M:=\operatorname{Spec}A\longrightarrow N$ be a finitely presented morphism to a derived Artin stack $N$. The relative de Rham algebra for $M\to N$ is the cdga $DR(M/N)\in\mathrm{cdga}^{\leq 0}_{A}$ given by
$DR(M/N):=\operatorname{Sym}_{A}(\mathbb{L}_{M/N}[1]))=\oplus^{\infty}_{p=0}\wedge^{p}\mathbb{L}_{M/N}[p],$ (3.29)
where $\mathbb{L}_{M/N}$ is the relative cotangent complex, and $DR(M/N)$ is equipped with the (cohomological) differential $d$ induced from the differential in $\mathbb{L}_{M/N}$.
We also have the de Rham differential $d_{DR}:\wedge^{p}\mathbb{L}_{M/N}\longrightarrow\wedge^{p+1}\mathbb{L}_{M/N}$. The de Rham differential defines a mixed structure on $DR(M/N)$ with $\epsilon:=d_{DR}$. This makes $DR(M/N)$ into a mixed cdga. In addition to the internal cohomological grading $DR(M/N)$ has a second (form degree) “weight grading” defined as $DR(M/N)(p):=\wedge^{p}\mathbb{\mathbb{missing}}L_{M/N}[p]$. The weight grading and mixed structure on $DR(M/N)$ are also compatible with the multiplicative structure and make $DR(M/N)$ into a graded mixed cdga over $\mathbb{k}$. The degree and weight of the internal differential $d$ are $1$ and 00, respectively. The degree and weight of $\epsilon$ are $-1$ and $1$, respectively.
Let $dAff_{/N}$ be the comma category of derived affine schemes mapping to $N$. The assignment $(\operatorname{Spec}A\longrightarrow N)\mapsto DR(\operatorname{Spec}A/N)$ defines a functor
$dAff_{/N}^{\text{op}}\longrightarrow\epsilon-dg^{gr}$
from $dAff_{/N}^{\text{op}}$ to the category of mixed graded cdga over $\mathbb{k}$. We can derive this functor on the left, by precomposing it with cofibrant replacement functor, to obtain a well-defined $\infty$-functor
$\mathbf{DR}(-/N):\mathbf{dAff}^{\text{op}}_{/N}\longrightarrow\epsilon-\mathbf{dg}^{gr}.$
Recall that for a mixed graded complex $F\in\epsilon-dg^{gr}$ of $\mathbb{k}$-vector spaces we have assoicated complexes $NC^{n}(F)(p):=\underset{i\geq 0}{\prod}F^{n-2i}(p+i)\in dg$ and a graded complex $NC^{w}(F):=\underset{p\in\mathbb{Z}}{\oplus}NC(F)(p)\in dg^{gr}$ of negative cyclic chains. Recall also that for any complex $E\in dg$ of $\mathbb{k}$-vector spaces we get a simplicial set $|E|=Map_{dg}(\mathbb{k},E)=$ the Dold-Kan functor applied to the $\tau_{\leq 0}$ truncation of $E$.
Definition 3.17.
For $A\in\mathrm{cdga}^{\leq 0}$ with a map $\operatorname{Spec}A\longrightarrow N$ of derived stacks and two integers $p\geq 0$ and $n\in\mathbb{Z}$, we define
  1. 1.
    The simplicial set $\mathcal{A}^{p}_{N}(\operatorname{Spec}(A),n)$ of relative $n$-shifted $p$-forms on $\operatorname{Spec}A/N$ by setting $\mathcal{A}^{p}_{N}(\operatorname{Spec}(A),n):=\left|\wedge^{p}\mathbb{L}_{\operatorname{Spec}(A)/N}[n]\right|$.
  2. 2.
    The simplicial set $\mathcal{A}^{p,cl}_{N}(\operatorname{Spec}(A),n)$ of relative closed $n$-shifted $p$-forms on $\operatorname{Spec}A/N$ by setting $\mathcal{A}^{p,cl}_{N}(\operatorname{Spec}(A),n):=|NC^{w}(DR(\operatorname{Spec}(A)/N))[n-p](p)|$.
The constructions $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ give rise to $\infty$-functors $\mathbf{dAff}^{op}_{/N}\longrightarrow\mathbb{S}$ from the $\infty$-category $\mathbf{dAff}^{op}_{/N}$ of affine derived schemes over $N$ to the $\infty$-category $\mathbb{S}$ of simplicial sets.
Definition 3.18.
For $\operatorname{Spec}(A)\in\mathbf{dAff}_{/N}$, the simplicial set $\mathcal{A}^{p}_{N}(\operatorname{Spec}(A),n)$ (respectively $\mathcal{A}^{p,cl}_{N}(\operatorname{Spec}(A),n)$) is called the space of $p$-forms of degree $n$ on the derived stack $\operatorname{Spec}(A)$, relative to $N$ (respectively the space of closed $p$-forms of degree $n$ on the derived stack $\operatorname{Spec}(A)$, relative to $N$).
The two $\infty$-functors $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived prestacks on $\mathbf{dAff}_{/N}$. By [32, Proposition 1.11], we know that $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived stacks for the big étale site over $N$.
Now let $S$ be an l.f.p. derived Artin stack which is equipped with a locally free log strurcture. Let $\mathbf{dSt}_{S,\log}$ be the $\infty$ category of derived Artin stacks $M$ equipped with a locally free log structure and a logarithmic map $M\to S$. In order to define relative shifted log $p$-forms for an object $M\longrightarrow S$ in $\mathbf{dSt}_{S,\log}$, we set $N:=\mathfrak{L}^{0}_{S,\mathsf{lf}}$ in the above construction. In other words we define stacks $\boldsymbol{\mathcal{A}}^{p}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)$, respectively $\boldsymbol{\mathcal{A}}^{p,cl}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)$ of relative log $p-$forms, respectively closed log $p-$ forms, of degree $n$ on $M\to S$ by setting $\boldsymbol{\mathcal{A}}^{p}_{S,\log}(-,n):=\boldsymbol{\mathcal{A}}^{p}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)$, respectively $\boldsymbol{\mathcal{A}}^{p,cl}_{S,\log}(-,n):=\boldsymbol{\mathcal{A}}^{p,cl}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)$.
Definition 3.19.
  1. (1)
    The space of $S$ relative logarithmic $n$-shifted $p$-forms on $M\in\mathbf{dSt}_{S,\log}$ is defined to be
    $\mathcal{A}_{log}^{p}(M/S,n):=\mathsf{Map}_{{}_{\textbf{dSt}_{S,log}}}\left(M,\boldsymbol{\mathcal{A}}^{p}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)\right).$
  2. (2)
    The space of closed relative logarithmic $n$-shifted $p$-forms on $M\in\mathbf{dSt}_{S,\log}$ is defined to be
    $\mathcal{A}_{log}^{p,cl}(M/S,n):=\mathsf{Map}_{{}_{\textbf{dSt}_{S,log}}}\left(M,\boldsymbol{\mathcal{A}}^{p,cl}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)\right).$
  3. (3)
    A $2$-form $\omega\in\mathcal{A}_{log}^{2}(M/S,n)$ is non-degenerate if the induced map in $\mathrm{D}_{qcoh}(M)$
    $\Theta_{\omega}:\mathbb{T}^{log}_{M/S}\longrightarrow\mathbb{L}^{log}_{M/S}[n]$ (3.30)
    is a quasi-isomorphism. We denote by $\mathcal{A}_{log}^{2}(M/S,n)^{nd}$ the union of all the connected components of $\mathcal{A}_{log}^{2}(M/S,n)$ which is consist of non-degenerate relative logarithmic $2$-forms of degree $n$ on $M$.
  4. (4)
    Finally, we define the space of $n$-shifted relative log symplectic forms as the homotopy fibre product
    $Symp_{log}(M/S,n):=\mathcal{A}_{log}^{2}(M/S,n)^{nd}\times^{h}_{\mathcal{A}_{log}^{2}(M/S,n)}\mathcal{A}_{log}^{2,cl}(M/S,n).$
4 The logarithmic Dolbeault moduli stack and its log symplectic structure
In this section, we will construct a logarithmic version of the relative Dolbeault moduli stack for the family of curves $\mathfrak{X}_{\mathfrak{M}}\rightarrow\mathfrak{M}$ defined at the end of section 3.2. We will show that the relative logarithmic Dolbeault moduli stack has a relative 00-shifted log-symplectic form over the spectrum of a discrete valuation ring $S$. Moreover, we will show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over the spectrum of a discrete valuation ring. This was proved for moduli schemes in [9].
Definition 4.1.
Let $\mathcal{X}\longrightarrow\mathcal{S}$ be a flat representable morphism of classical Artin stacks of relative dimension $1$. We call it a semi-stable family of curves if every geometric fibre of $\mathcal{X}\longrightarrow\mathcal{S}$ is isomorphic to a semi-stable curve i.e., it satisfies the following properties
  1. 1.
    every geometric fibre is a reduced, connected, projective, nodal curve
  2. 2.
    every component $E$ of a geometric fiber which isomorphic to $\mathbb{P}^{1}$ must intersect the union of all other components at at least two smooth points.
Recall that if $C$ is a nodal curve with a set of nodes $D$ we have an explicit formula for the dualizing sheaf of $C$. Indeed, let $q:\widetilde{C}\longrightarrow C$ denote the normalisation and let $\widetilde{D}$ denote the pre-image $q^{-1}(D)$. Then the dualizing sheaf $\omega_{C}$ of $C$ is the kernel
$\omega_{C}=\ker\left[q_{*}\Omega_{\widetilde{C}}^{1}(\widetilde{D})\longrightarrow\underset{x\in D}{\bigoplus}\ \mathbb{k}_{x}\right],$ (4.1)
where
  • $\mathbb{k}_{x}$ denotes the sky-scraper sheaf at the point $x$.
  • for the two preimages $q^{-1}(x)=\{x^{-},x^{+}\}$ of a node $x\in D$, the map $q_{*}\Omega_{\widetilde{C}}^{1}(x^{+}+x^{-})\longrightarrow\mathbb{k}_{x}$ is given by
    $s\mapsto\mathsf{Res}(s;x^{+})+\mathsf{Res}(s;x^{-}),$ (4.2)
    where $\mathsf{Res}(s;x)$ denotes the residue of a form $s$ at a point $x$.
It is straightforward to check that it is a locally free sheaf of rank one. It is well known that this property persists in families. That is, for a family $f:\mathcal{X}\to\mathcal{S}$ of semi-stable/ pre-stable curves there is a well defined relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$ which is a line bundle on the total space of the family [39, 109.19 The relative dualizing sheaf]. In the case when $\mathcal{S}$ and $\mathcal{X}$ are both smooth, the relative dualizing sheaf is naturally identified with the sheaf of relative logarithmic forms along the fibers of $f$. Concretely, if $\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\subset\mathcal{S}$ denotes the discriminant divisor of the map $f:\mathcal{X}\to\mathcal{S}$, and $\boldsymbol{\mathfrak{d}}_{\mathcal{X}}=f^{-1}\left(\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\right)$ is the normal crossings divisor in $\mathcal{X}$ comprised of the singular fibers of $f$, then we have
$\omega_{\mathcal{X}/\mathcal{S}}=\Omega^{1}_{\mathcal{X}}(\log\boldsymbol{\mathfrak{d}}_{\mathcal{X}})/f^{*}\Omega^{1}_{\mathcal{X}}(\log\boldsymbol{\mathfrak{d}}_{\mathcal{S}}).$
Alternatively, again under the assumption that $\mathcal{X}$ and $\mathcal{S}$ are smooth, we can identify the relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$ with the relative cotangent sheaf $\Omega^{1}_{l\mathcal{X}/l\mathcal{S}}$ of the log smooth map $l\mathcal{X}\to l\mathcal{S}$, where $l\mathcal{X}$ and $l\mathcal{S}$ are the log stacks whose underlying stacks are $\mathcal{X}$ and $\mathcal{S}$ and whose log structures are given by the normal crossings divisors $\boldsymbol{\mathfrak{d}}_{\mathcal{X}}$ and $\boldsymbol{\mathfrak{d}}_{\mathcal{S}}$ respectively. More generally, for any semi-stable family of curves $\mathcal{X}\to\mathcal{S}$ for which $\mathcal{S}$ is a classical Artin stack which is locally of finite type, the construction of F. Kato [21] implies that there are canonical functorial log structures $l\mathcal{X}$ and $l\mathcal{S}$ for which the map $\mathcal{X}\to\mathcal{S}$ can be refined to a log smooth morphism $l\mathcal{X}\to l\mathcal{S}$ so that still have $\omega_{\mathcal{X}/\mathcal{S}}\cong\Omega^{1}_{l\mathcal{X}/l\mathcal{S}}$.
Definition 4.2.
A relative Higgs bundle over a family of semistable curves $\mathcal{X}/\mathcal{S}$ is a locally free $\mathcal{O}_{\mathcal{X}}-$module $\mathcal{E}$ with a Higgs field $\phi:\mathcal{E}\longrightarrow\mathcal{E}\otimes\omega_{\mathcal{X}/\mathcal{S}}$.
Remark 4.3.
The vector bundle $\mathcal{E}$ along with the Higgs field $\phi:\mathcal{E}\to\mathcal{E}\otimes\omega_{\mathcal{X}/\mathcal{S}}$ might also be dubbed a relative logarithmic Higgs bundle on $\mathcal{X}/\mathcal{S}$ since, as explained above, such a $\phi$ take values in $\Omega^{1}_{l\mathcal{X}/l\mathcal{S}}$. In the previous definition, we chose to call such a pair $(\mathcal{E},\phi)$ simply a relative Higgs bundle since it is just a Higgs bundle on $\mathcal{X}$ with coefficients in the relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$ of the map $\mathcal{X}\to\mathcal{S}$.
Next we consider $\mathsf{T}_{\mathcal{X}/\mathcal{S}}:=\underline{Spec}_{\mathcal{X}}Sym_{\mathcal{O}_{\mathcal{X}}}(\omega_{\mathcal{X}/\mathcal{S}})$. This is the total space of the line bundle dual to the relative dualizing sheaf of the family of curves $\mathcal{X}\longrightarrow\mathcal{S}$.
Definition 4.4.
Let $\widehat{\mathsf{T}_{\mathcal{X}/\mathcal{S}}}$ denote the formal completion of $\mathsf{T}_{\mathcal{X}/\mathcal{S}}$ along the zero section. Notice that it is a formal group scheme over $\mathcal{X}$, whereas $\mathsf{T}_{\mathcal{X}/\mathcal{S}}$ is an abelian group scheme over $\mathcal{X}$.
  • The (logarithmic) Dolbeault stack of $\mathcal{X}/\mathcal{S}$ is defined as the formal quotient stack
    $\mathcal{X}_{Dol/\mathcal{S}}:=\left[\widehat{\mathsf{T}_{\mathcal{X}/\mathcal{S}}}\times_{S}\mathcal{X}\rightrightarrows\mathcal{X}\right]$ (4.3)
  • The nilpotent (logarithmic) Dolbeault stack of $\mathcal{X}/\mathcal{S}$ is defined as the quotient stack
    $\mathcal{X}^{nil}_{Dol/\mathcal{S}}:=[\mathsf{T}_{\mathcal{X}/\mathcal{S}}\times_{\mathcal{S}}\mathcal{X}\rightrightarrows\mathcal{X}]$ (4.4)
With this notation, we have the following standard lemma
Lemma 4.5.
  1. 1.
    $\mathrm{QCoh}(\mathcal{X}_{Dol/\mathcal{S}})\cong\mathrm{Mod}_{\operatorname{Sym}_{\mathcal{O}_{\mathcal{X}}}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})}(\mathrm{QCoh}(\mathcal{X}))$
  2. 2.
    If $(\mathcal{E},\phi)$ is a quasi-coherent relative Higgs complex on $\mathcal{X}/\mathcal{S}$ and if $E$ is the corresponding quasi-coherent complex on $\mathcal{X}_{Dol/\mathcal{S}}$ then we have an isomorphism of quasi-coherent complexes on $\mathcal{S}$:
[…]
Proof.
The proof follows verbatim [31, Proposition 5.1.2] after replacing the relative cotangent complex $\mathbb{L}_{\mathcal{X}/\mathcal{S}}$ with the relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$. ∎
Remark 4.6.
Notice that $\mathcal{X}_{Dol/\mathcal{S}}$ is the relative classifying stack $B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}$ of the formal commutative group scheme $\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}$ over $\mathcal{X}$. The nilpotent Dolbeault stack $\mathcal{X}^{nil}_{Dol/\mathcal{S}}$ is the classifying stack of the abelian group scheme $\mathsf{T}_{\mathcal{X}/\mathcal{S}}$ over $\mathcal{X}$. Therefore we have the following equivalence of categories.
{Quasi-coherent $\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves}\\ \text{over}\leavevmode\nobreak\ B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \text{with}\\ \text{an action of the sheaf of algebras}\\ \operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}$ (4.5)
Notice that $\operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})$ is a quadratic algebra. The quadratic dual algebra $(\operatorname{Sym}\leavevmode\nobreak\ \omega^{\vee}_{\mathcal{X}/\mathcal{S}})^{!}$ is isomorphic to the dg-algebra $\operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1])$ with zero differential. Moreover, we have the following equivalence of categories ([30]).
{Quasi-coherent sheaves $\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \text{with}\\ \text{an action of the sheaf of algebras}\\ \operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\\ \operatorname{Spec}_{\mathcal{X}}\leavevmode\nobreak\ \operatorname{Sym}\leavevmode\nobreak\ (\omega_{\mathcal{X}/\mathcal{S}}[-1])\end{array}\right\}$ (4.6)
From (4.5) and (4.6), we have the following.
$B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\cong\operatorname{Spec}_{\mathcal{X}}\leavevmode\nobreak\ \operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1]).$ (4.7)
Similarly, we have the following equivalence.
{Vector $\left\{\begin{array}[]{@{}ll@{}}\text{Vector bundles}\\ \text{over}\leavevmode\nobreak\ B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Vector bundles over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{with an}\\ \text{ action of the sheaf of algebras}\\ ({\operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}}))^{\vee}\cong\widehat{\operatorname{Sym}}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{nilpotent Higgs}\\ \text{bundles over}\leavevmode\nobreak\ \mathcal{X}\end{array}\right\}$ (4.8)
Theorem 4.7.
$\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $\mathcal{S}$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 00-shifted relative symplectic structure over $\mathcal{S}$.
Proof.
Proof of $\mathcal{O}$-compactness: Let us begin with recalling the notations for the following maps from Section 1.1.
${\mathcal{X}_{Dol/\mathcal{S}}}$${\mathcal{X}}$${\mathcal{S}}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$ (4.9)
From (4.7), we have
$Rq_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\cong\mathcal{O}_{\mathcal{X}}\oplus\omega_{\mathcal{X}/\mathcal{S}}[-1],$ (4.10)
which is clearly a perfect complex. Therefore, $\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}$ is $\mathcal{O}$-compact relative to $\mathcal{S}$.
Proof of $\mathcal{O}$-orientation: Using (4.10), we have
$(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}})^{\vee}[-2]$ $\cong(Rr_{*}\mathcal{O}_{\mathcal{X}}\oplus Rr_{*}\omega_{\mathcal{X}/\mathcal{S}}[-1])^{\vee}[-2]$
$\cong(Rr_{*}\mathcal{O}_{\mathcal{X}})^{\vee}[-2]\oplus(Rr_{*}\omega_{\mathcal{X}/\mathcal{S}}[-1])^{\vee}[-2].$
By Serre duality, we have
$Rr_{*}(\omega_{\mathcal{X}/\mathcal{S}}[-1])\cong(Rr_{*}((\omega_{\mathcal{X}/\mathcal{S}}[-1])^{\vee}\otimes\omega_{\mathcal{X}/\mathcal{S}}[1]))^{\vee}\cong(Rr_{*}(\mathcal{O}_{\mathcal{X}}[2]))^{\vee}\cong(Rr_{*}\mathcal{O}_{\mathcal{X}})^{\vee}[-2]$
$\implies(Rr_{*}(\omega_{\mathcal{X}/\mathcal{S}}[-1]))^{\vee}[-2]\cong Rr_{*}(\mathcal{O}_{\mathcal{X}}[2])[-2]\cong Rr_{*}\mathcal{O}_{\mathcal{X}}\hskip 60.00009pt$
Notice that $H^{0}(Rr_{*}\mathcal{O}_{\mathcal{X}})\cong\mathcal{O}_{\mathcal{S}}$. We choose an isomorphism once and for all and denote it by $\eta$. This defines an element
$\eta\in(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}})^{\vee}[-2]=R\mathcal{H}om_{\mathcal{S}}(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}},\mathcal{O}_{\mathcal{S}}[-2]).$ (4.11)
We want to show that it is an $\mathcal{O}-$orientation.
Let $\mathcal{E}$ be a perfect complex on $\mathcal{X}_{Dol/\mathcal{S}}\times_{\mathcal{S}}\textbf{Spec}(A)$, for an $A\in\textbf{cdga}^{\leq 0}_{{\mathcal{S}}}$. We have to show that the following map
$(-\cap\eta_{A}):R{p_{A}}_{*}\mathcal{E}\longrightarrow(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}[-2]=(R{p_{A}}_{*}(\mathcal{E}^{\vee})[2])^{\vee}$ (4.12)
is a quasi-isomorphism of $A$-dg-modules, where $\eta_{A}:Rp_{A*}\mathcal{O}_{\mathcal{X}_{Dol,A}}\longrightarrow A[-2]$ is the derived pullback of $\eta:Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\longrightarrow\mathcal{O}_{\mathcal{S}}[-2]$ under the map ${\rm Spec}\,(A)\longrightarrow\mathcal{S}$ and $(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}$ is the derived $A$-dual of $R{p_{A}}_{*}(\mathcal{E}^{\vee})$. Here
$p_{A}:\mathcal{X}_{Dol/\mathcal{S}}\times_{\mathcal{S}}\textbf{Spec}(A)\longrightarrow\textbf{Spec}(A)$
is the projection. The left hand side is isomorphic to $R{r_{A}}_{*}\mathsf{Dol}_{(E_{A},\phi_{A})}$, where
$\mathsf{Dol}_{(E_{A},\phi_{A})}:=[E_{A}\xrightarrow{\phi_{A}}E_{A}\otimes\omega_{\mathcal{X}_{A},A}]$
$0\hskip 113.81102pt1$
corresponding to $R{q_{A}}_{*}\mathcal{E}_{A}$ (here, $E_{A}$ is sitting in degree 00). The right hand side is isomorphic to $(R{r_{A}}_{*}(\mathsf{Dol}^{\mathsf{D}}_{(E_{A},\phi_{A})}))^{\vee}$, where
$\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}:=[E^{\vee}_{A}\xrightarrow{-\phi^{\vee}_{A}}E^{\vee}_{A}\otimes\omega_{\mathcal{X}_{A},A}]$ (4.13)
$-2\hskip 99.58464pt-1$ (4.14)
where $\phi^{\vee}_{A}$ is the image of $\phi_{A}$ under the natural isomorphism $\mathcal{E}nd(E_{A})\cong\mathcal{E}nd(E^{\vee}_{A})$. Now since the Grothendieck-Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$ is the same as $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$, the morphism (4.12) is quasi-isomorphism. (We have used the notation $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$ in (4.13) to disambiguate between the ordinary dual and the Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$.)
Therefore, we have shown that $\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over ${\mathcal{S}}$ and from [32, Theorem 2.5], it follows that $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 00-shifted relative symplectic structure over ${\mathcal{S}}$. ∎
Let us fix an integer $n$, which will represent rank in the moduli problem. Let $S$ be a spectrum of a discrete valuation ring over $\mathbb{k}$. We start with the set-up as in 3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
Proposition 4.8.
The morphism $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))\longrightarrow\mathfrak{M}$ is a quasi-smooth morphism of derived Artin stacks.
Proof.
Let us denote $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))$ by $M$, for simplicity of notation. As before, the natural map $p:\mathfrak{X}_{Dol,\mathfrak{M}}\to\mathfrak{M}$ factors as
${\mathfrak{X}_{Dol/\mathfrak{M}}}$${\mathfrak{X}}$${\mathfrak{M}}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$ (4.15)
The relative tangent complex of the morphism $M\longrightarrow\mathfrak{M}$ is given by
$\mathbb{T}_{M/\mathfrak{M}}\cong Rp_{*}\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}[1]\cong Rr_{*}\mathcal{C}(E,\phi)[1]$ (4.16)
where $\mathcal{C}(E,\phi)$ denotes the following complex
$\left[\mathcal{E}nd(E)\xrightarrow{[-,\phi]}\mathcal{E}nd(E)\otimes\mathsf{pr}_{\mathfrak{X}}^{*}\omega_{\mathfrak{X}/\mathfrak{M}}\right]$ (4.17)
with $\mathcal{E}nd(E)$ sitting in the degree 00. Here $\mathcal{E}$ denotes the universal sheaf on $\mathfrak{X}_{Dol/\mathfrak{M}}\times_{\mathfrak{M}}M$ and $(E,\phi)$ denotes the corresponding universal relative Higgs complex on $\mathfrak{X}\times_{\mathfrak{M}}M$. From [40, Theorem 0.3], it follows that $\mathbb{L}_{M/\mathfrak{M}}$ is a perfect complex because the map $r:\mathfrak{X}\longrightarrow\mathfrak{M}$ is a proper representable local complete intersection morphism. Also, using Grothendieck-Serre duality we can see that the complex $Rr_{*}\mathcal{C}(E,\phi)[1]$ has amplitude in $[-1,1]$. Therefore, we conclude that the morphism
$\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))\longrightarrow\mathfrak{M}$
is a quasi-smooth morphism of derived Artin stacks. ∎
Theorem 4.9.
The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))$ has a natural relative 00-shifted log-symplectic structure over $S$.
Proof.
Recall that $\mathfrak{L}^{0}_{S,\mathsf{lf}}$ denotes the classifying stack of log morphisms to $S$ from schemes equipped with locally free log structures. The log structure on $\mathfrak{M}$, which we discussed in Proposition 3.11, then gives a map $\mathfrak{M}\to\mathfrak{L}^{0}_{S,\mathsf{lf}}$.
Consider the composite morphism.
$M:=\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{L}^{0}_{S,\mathsf{lf}}$
The composite morphism induces a log structure on $M$, which is the same as the pull-back of the log structure from $\mathfrak{M}$. Notice that the map $M\to\mathfrak{M}$ is quasi-smooth (Proposition 4.8) and the map $\mathfrak{M}\to\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is smooth with $\mathbb{L}_{\mathfrak{M}/\mathfrak{L}^{0}_{S,\mathsf{lf}}}\cong 0$ (Proposition 3.11). Therefore, it follows that
$\mathbb{L}_{M/\mathfrak{L}^{0}_{S,\mathsf{lf}}}\cong\mathbb{L}_{M/\mathfrak{M}}\cong Rr_{*}\mathcal{C}(E,\phi)[1].$ (4.18)
So, the relative log-cotangent complex of the map $M\longrightarrow S$ is isomorphic to the relative cotangent complex of the map $M\longrightarrow\mathfrak{M}$. From 4.7, it follows that the stack $M$ has a 00-shifted symplectic structure relative to the stack $\mathfrak{M}$, which is, by definition 3.19, a 00-shifted log-symplectic structure on $M$ relative to $S$.
The log-symplectic pairing can be described as follows. The stack $BGL_{n}$ has a $2$-shifted symplectic form given by the trace pairing:
$\mathbb{T}_{BGL_{n}}\wedge\mathbb{T}_{BGL_{n}}=\mathfrak{g}l_{n}[1]\wedge\mathfrak{g}l_{n}[1]\xrightarrow{(\mathsf{x},\mathsf{y})\mapsto Trace(\mathsf{x}\cdot\mathsf{y})}\mathbb{k}[2].$ (4.19)
Now let $A$ be an object of $\mathrm{cdga}^{\leq 0}$ and let $\operatorname{Spec}(A)\to M$ be an $A$-valued point of $M$. Write $\mathfrak{X}_{A}$ for the pullback $\mathfrak{X}\times_{\mathfrak{M}}\operatorname{Spec}(A)$ of the universal curve to $\operatorname{Spec}(A)\to M\to\mathfrak{M}$. We will also write $\mathfrak{X}_{Dol,A}$ for the pullback $\mathfrak{X}_{Dol/\mathfrak{M}}\times_{\mathfrak{M}}\operatorname{Spec}(A)$ for the pullback of the relative Dolbeault stack. Now, by the definition of $M$, the map $\operatorname{Spec}(A)\to M$ gives us a map
$f_{A}:\mathfrak{X}_{Dol,A}\longrightarrow BGL_{n}\times\operatorname{Spec}(A)$
which corresponds to a vector bundle $\mathcal{E}_{A}$ on $\mathfrak{X}_{Dol,A}$. The vector bundle $\mathcal{E}_{A}$ on $\mathfrak{X}_{Dol,A}$ corresponds to a Higgs field
$(E_{A}\xrightarrow{\phi_{A}}E_{A}\otimes\omega_{\mathfrak{X}_{A},A}).$
on $\mathfrak{X}_{A}$.
We have the following induced pairing by pulling back (4.19).
$f^{*}_{A}\mathbb{T}_{BGL_{n}}\otimes f^{*}_{A}\mathbb{T}_{BGL_{n}}=\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\xrightarrow{Tr:=Trace}\mathcal{O}_{\mathfrak{X}_{Dol,A}}[2]$ (4.20)
Now by pushing it forward by the map $q_{A}$, we get
${{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ E_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ E_{A}[1])}$${(\mathcal{O}_{\mathcal{X}_{A}}\oplus\omega_{\mathfrak{X}_{A},A}[-1])[2]}$ (4.21)
The complex ${q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])=[\mathcal{E}nd\leavevmode\nobreak\ E_{A}\xrightarrow{[-,\phi_{A}]}\mathcal{E}nd\leavevmode\nobreak\ E_{A}\otimes\omega_{\mathfrak{X}_{A},A}]$ can be used to describe the pullback of the pairing in (4.21) by the natural map
${q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])\otimes{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])\to{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])$
as the following map of complexes.
[…]
where the map $d_{1}:=([-,\phi],[-,\phi])$ and $d_{2}:=[-,\phi]\otimes\mathbb{1}+\mathbb{1}\otimes[-,\phi]$. Finally, by pushing forward (4.21) by the map $r_{A}$ and then composing with the orientation $\eta_{A}$ (see (4.12)), we get our log-symplectic pairing. From the above description it is evident that this log-symplectic form coincides with the log-symplectic form discussed in [9]. ∎
5 Completeness of the Hitchin map
In this section, we define the Hitchin map on the ordinary Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see 8.1.2). We prove that the Hitchin map is complete. This result was proved in [1] for the Hitchin map as defined on the moduli scheme of stable Higgs bundles in the case where rank and degree are co-prime. We prove it here for the moduli stack, and the argument does not require us to assume that the rank and degree are co-prime.
Definition 5.1.
[38, p7] Working over $\operatorname{Spec}\mathbb{k}$, a morphism of Artin stacks $f:\mathcal{Y}_{1}\longrightarrow\mathcal{Y}_{2}$ is called complete if given any discrete valuation ring $A$ over $\mathbb{k}$ with fraction field $K$ and any commutative diagram
${\operatorname{Spec}\leavevmode\nobreak\ K}$${\mathcal{Y}_{1}}$${\operatorname{Spec}\leavevmode\nobreak\ A}$${\mathcal{Y}_{2}}$$\scriptstyle{f}$ (5.1)
there exists finite field extension $K\longrightarrow K^{\prime}$ with $A^{\prime}$ the integral closure of $A$ in $K^{\prime}$ and a lift over (the dotted arrow) $\operatorname{Spec}\leavevmode\nobreak\ A^{\prime}\longrightarrow\mathcal{X}$ making all the triangles in the diagram commute.
${\operatorname{Spec}\leavevmode\nobreak\ K^{\prime}}$${\operatorname{Spec}\leavevmode\nobreak\ K}$${\mathcal{Y}_{1}}$${\operatorname{Spec}\leavevmode\nobreak\ A^{\prime}}$${\operatorname{Spec}\leavevmode\nobreak\ A}$${\mathcal{Y}_{2}}$$\scriptstyle{f}$ (5.2)
A morphism of derived Artin stacks is called complete if the underlying morphism of ordinary Artin stacks is complete.
We recall the following setup from before. Let $S$ be a spectrum of a discrete valuation ring over $\mathbb{k}$. We start with the set-up as in 3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
Definition 5.2.
(Hitchin map) We recall the following two well-known definitions for the family of curves $\pi:\mathcal{X}\to S.$
  1. 1.
    $B:=\mathrm{Tot}(\oplus^{n}_{i=1}R^{0}\pi_{*}\omega_{\mathcal{X}/S})$. It is a vector bundle over the spectrum $S$ of a discrete valuation ring over $\mathbb{k}$. As is standard, we call $B$ the Hitchin base.
  2. 2.
    There is a natural map
    $h:M:=\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))\longrightarrow B,$
    which sends a family of Higgs bundles $(\mathfrak{X}_{T},\mathcal{E}_{T},\phi_{T})$ parametrised by an affine scheme $T$ over $S$ to
    $(-Trace(\phi_{T}),Trace(\wedge^{2}\phi_{T}),\dots,(-1)^{i}Trace(\wedge^{i}\phi_{T}),\dots,(-1)^{n}Trace(\wedge^{n}\phi_{T})).$
    inside $T\times_{S}B$
Now consider the substack $M^{cl}_{Gie}\subseteq\tau_{0}(M)\subseteq M$ in the classical truncation $\tau_{0}(M)$ of $M$ parametrizing Higgs bundles whose underlying vector bundle satisfies the Gieseker conditions. Namely, it is globally generated on the fibers of the map $X_{k}\to X_{0}$ and whose pushforward to $X_{0}$ is torsion-free. See 8.1.2 for a detailed description. Note also that for semi-stable families $\mathfrak{X}/S$ of arithmetic genus greater than or equal to $2$, the morphism $\tau_{0}(M)\to S$ is l.c.i. and therefore $\tau_{0}(M)\cong M$. In particular, the derived structure on $M$ plays a role only in genus zero and one.
The Hitchin map $h$ (see definition 5.2) defines a map $M^{cl}_{Gie}\to B$ by restriction.
Theorem 5.3.
The morphism $h|_{M^{cl}_{Gie}}:M^{cl}_{Gie}\longrightarrow B$ is complete.
Proof.
Let $T:=\operatorname{Spec}A$ be a spectrum of a discrete valuation ring over $\mathbb{k}$ with function field $K$. Let us denote $T^{o}:=\operatorname{Spec}K$. Suppose we are given a commutative diagram
[…]
We need to prove that this diagram factorizes through a map $T\to M^{cl}_{Gie}$. Let $TFH(\mathcal{X}/S)$ denote the ordinary Artin stack of families of torsion-free $\omega_{\mathcal{X}/S}$-valued Higgs pairs on the family of curves $\mathcal{X}/S$ (see Definition 8.1.1). Then there is a morphism
$\theta:M^{cl}_{Gie}\rightarrow TFH(\mathcal{X}/S)$
$(\pi:\mathcal{X}^{mod}\to\mathcal{X},\mathcal{E},\phi)\to(\pi_{*}\mathcal{E},\pi_{*}\phi)$
given by the pushforward of Higgs bundles under some modification $\mathcal{X}^{\textsf{mod}}\rightarrow\mathcal{X}$ of Gieseker curves, as defined in 3.2. Therefore, from (5.3) we get the following diagram
[…]
Using [15, Proposition 6] and [24, Lemma 6.5], we have a map $T\to TFH(\mathcal{X}/S)$ making the triangle commute:
[…]
Notice now:
  1. 1.
    if the spectrum $T$ of the discrete valuation ring is not faithfully flat over $S$ then the surface $\mathcal{X}\times_{S}T$ is isomorphic to the product $X_{0}\times T$
  2. 2.
    if the spectrum $T$ of the discrete valuation ring is faithfully flat over $S$ then the surface $\mathcal{X}\times_{S}T$ is isomorphic to a normal surface with an isolated singularity of type $\frac{\mathbb{k}[[x,y,t]]}{xy-t^{d}}$.
In the second case, it is a routine check that [15, Proposition 6] holds for the normal surface. Indeed, any vector bundle $\mathcal{E}^{o}$ on the generic fibre $\mathcal{X}\times_{S}T^{o}=\mathcal{X}\times_{S}\operatorname{Spec}(K)$ can be extended to the surface $\mathcal{X}\times_{S}T$ as a torsion-free sheaf $\mathcal{F}$. Write $i:\mathcal{X}\times_{S}T^{o}\hookrightarrow\mathcal{X}\times_{S}T$ for the natural inclusion, and suppose we are given a Higgs field $\phi^{o}:\mathcal{E}^{o}\to\mathcal{E}^{o}\otimes i^{*}\omega_{\mathcal{X}_{T}/T}$ on $\mathcal{E}^{o}$. To extend $\phi^{o}$ over all of $\mathcal{X}\times_{S}T$ consider the following diagram
${\mathcal{F}}$${i_{*}\mathcal{E}^{o}}$${i_{*}\mathcal{E}^{o}/{\mathcal{F}}}$00{0}00{0}${\mathcal{F}\otimes\omega_{\mathcal{X}_{T}/T}}$${i_{*}(\mathcal{E}^{o}\otimes i^{*}\omega_{\mathcal{X}_{T}/T})}$${(i_{*}\mathcal{E}^{o}/{\mathcal{F}})\otimes\omega_{\mathcal{X}_{T}/T}}$00{0}$\scriptstyle{i_{*}\phi^{o}}$ (5.6)
We wish to prove the existence of a Higgs field $\mathcal{F}\to\mathcal{F}\otimes\omega_{\mathcal{X}_{T}/T}$ making the diagram commute. Using [24, Lemma 6.5], we can extend the Higgs field on $\mathcal{F}$ except at the node. Therefore, the composite map $\mathcal{F}\longrightarrow(i_{*}\mathcal{E}^{o}/\mathcal{F})\otimes\omega_{\mathcal{X}_{T}/T}$ vanishes at the generic point of the closed fibre $X_{0}$. Therefore, the obstruction to the existence of the Higgs field we seek lives in the torsion subsheaf of $i_{*}\mathcal{E}^{o}/\mathcal{F}$, when $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is viewed as a sheaf on its support $X_{0}$. Therefore, if we show that the sheaf $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is pure, i.e. is torsion-free on its support $X_{0}$, then we will show that the induced Higgs field on $\mathcal{F}$ extends everywhere. To see this we restrict the upper exact sequence to the closed fibre $X_{0}$. We get the exact sequence
$0\longrightarrow\mathcal{T}or^{1}_{{}_{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})\longrightarrow\mathcal{F}|_{X_{0}}\longrightarrow(i_{*}\mathcal{E}^{o})|_{X_{0}}\longrightarrow i_{*}\mathcal{E}^{o}/{\mathcal{F}}\longrightarrow 0$ (5.7)
However, $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})\cong(i_{*}\mathcal{E}^{o}/{\mathcal{F}})\otimes\mathcal{O}_{{}_{\mathcal{X}_{T}}}(-X_{0})$. Since $\mathcal{O}_{{}_{\mathcal{X}_{T}}}(-X_{0})$ is a locally-free sheaf, $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is pure if and only if $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})$ is pure. But $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})$, viewed as a sheaf on its support, is a subsheaf of $\mathcal{F}|_{X_{0}}$, therefore as a sheaf on $X_{0}$ it is torsion-free. This completes the argument in the second case.
In the first case, the surface $\mathcal{X}_{T}\cong X_{0}\times T$. Therefore, the normalisation of $\mathcal{X}_{T}$ is isomorphic to $\widetilde{X_{0}}\times T$. Let us denote the normalisation morphism by $f:\widetilde{X_{0}}\to X_{0}$. From [10, Proposition 3.2], it follows that we can extend the pullback of the generic Higgs bundle $((f\times\mathrm{id}_{T^{o}})^{*}\mathcal{E}^{o},(f\times\mathrm{id}_{T^{o}})^{*}\phi^{o})$ as a good generalized parabolic Hitchin pair which in turn gives us the desired torsion-free Hitchin pair on the surface $\mathcal{X}_{T}$.
So now, since we have extended the generic Higgs bundle as a torsion-free Hitchin pair, we obtain the following commutative diagram
[…]
We now choose a suitable connected component $R^{\Lambda,m_{T}}_{S}$ of the atlas of $TFH(\mathcal{X}/S)$, which was constructed in 8.1.1. The component is chosen so that the image of the map $T\longrightarrow TFH(\mathcal{X}/S)$ lies inside the image of the map $R^{\Lambda,m_{T}}_{S}\longrightarrow TFH(\mathcal{X}/S)$. Let $\mathcal{Y}^{H}_{S}$ denote the quot scheme for Gieseker-Hitchin pairs constructed over the quot scheme $R^{\Lambda,m_{T}}_{S}$ (see 8.1.4). Let us denote the projection map by $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$. Notice that the maps $R^{\Lambda,m_{T}}_{S}\longrightarrow TFH(\mathcal{X}/S)$ and $\mathcal{Y}^{H}_{S}\longrightarrow M^{cl}_{Gie}$ are principal $GL_{N}$-bundles for some large integer $N$ and the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is a $GL_{N}$-equivariant morphism (see 8.8).
[…]
Since $T$ is a spectrum of a discrete valuation ring, any $GL_{N}$ bundle on $T$ is trivial. Therefore, by choosing a trivialisation of the $GL_{N}$-bundle $R^{\Lambda,m_{T}}_{S}\times_{TFH(\mathcal{X}/S)}T\to T$ we get the following lift of the diagram 5.8.
[…]
From [1, Proposition 5.11], it follows that the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper (We discuss the proof below in Lemma 5.4 and Corollary 5.5 for reader’s convenience). So there exists a lift $T\longrightarrow\mathcal{Y}^{H}_{S}$. The composite map $T\longrightarrow\mathcal{Y}^{H}_{S}\longrightarrow M^{cl}_{Gie}$ is our desired extension. ∎
Lemma 5.4.
Let $Y\longrightarrow R$ be a quasi-projective morphism of separated schemes over a spectrum of a discrete valuation ring $S$ which is an isomorphism over the generic point of $S$. Suppose any map $T\longrightarrow R$ flat over $S$ can be lifted to a map $T\longrightarrow Y$. Then the map $Y\longrightarrow R$ is proper.
Proof.
Take the closure of $Y$ inside some projective bundle (over $R$) to get a projective morphism $\overline{Y}\longrightarrow R$. Take any element $y\in\overline{Y}\setminus Y$. Then there exists a map $T\longrightarrow\overline{Y}$ (here, $T$ is a spectrum of a d.v.r) passing through $y$ and flat over $S$. Therefore we get a map $T\longrightarrow R$. But then, by our lifting assumption, this map can be lifted to a map $T\longrightarrow Y$. Let $y^{\prime}\in Y$ be the image of the closed point of $T$ under this map. But, by separateness of the $\overline{Y}$, we must have $y=y^{\prime}$ and therefore $\overline{Y}=Y$, and the map $Y\longrightarrow R$ is proper. ∎
Corollary 5.5.
[1, Proposition 5.11] The map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper.
Proof.
Since the map $\tilde{\theta}$ is quasi-porjective, we choose a closure $\overline{\mathcal{Y}^{H}_{S}}\longrightarrow R^{\Lambda,m_{T}}_{S}$ inside some projective bundle over $R^{\Lambda,m_{T}}_{S}$. Let $y\in\overline{\mathcal{Y}^{H}_{S}}\setminus{\mathcal{Y}^{H}_{S}}$. Then there exists a map $T\longrightarrow\overline{\mathcal{Y}^{H}_{S}}$ (here, $T$ is a spectrum of a d.v.r) passing through $y$ and flat over $S$. Consider the composite map $T\to\overline{\mathcal{Y}^{H}_{S}}\to R^{\Lambda,m_{T}}_{S}$. Let $(\mathcal{F}_{T},\phi_{T})$ be the flat family of torsion-free Higgs pair on $\mathcal{X}_{T}:=\mathcal{X}\times_{S}T$ corresponding to the map $T\to R^{\Lambda,m_{T}}_{S}$.
Claim: The map $T\to R^{\Lambda,m_{T}}_{S}$ can be lifted to a map $T\to\mathcal{Y}^{H}_{S}$.
Proof of the claim: Notice that the surface $\mathcal{X}_{T}$ may have a singularity of type $\frac{\mathbb{k}[|x,y,t|]}{(xy-t^{n})}$ because the map $T\to S$ is a faithfully-flat map of spectrum of discrete valuation rings. Let $r_{{}_{T}}:\mathcal{X}^{{res}}_{T}\longrightarrow\mathcal{X}_{T}$ be the minimal resolution of singularities. Then from [18, Proposition 6.5] and [1, §4] it follows that the vector bundle $\mathcal{E}_{T}:=\frac{r_{{}_{T}}^{*}\mathcal{F}_{T}}{Torsion}$ has the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$. Notice also that by construction and by the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$, it follows that the natural map $r_{{}_{T}}^{*}(r_{{}_{T}})_{*}\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}$ is a surjective map of sheaves, which means $(\mathcal{E}_{T})|_{R}$ is globally generated. Here $R$ denotes the chain of rational curves in $\mathcal{X}^{{res}}_{T}$. Hence, by definition (Definition 8.1 and Remark 8.2), the vector bundle $\mathcal{E}_{T}$ is a Gieseker vector bundle. The pullback of the Higgs field $\phi_{{}_{T}}$ defines a Higgs field on $\mathcal{E}_{T}$. The pair $(\mathcal{E}_{T},\phi_{T})$ is the desired lift.
Therefore from Lemma 5.4 it follows that the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper.
6 Flatness of the Hitchin map
In this section, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$ (see 8.1.2), which is the reduced fibre of $h$ over the point $0\in B$. We prove that every irreducible component of of the reduced nilpotent cone has an open subset which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat.
Definition 6.1.
The nilpotent cone is the Hitchin fibre over the zero section $0_{S}$ of $B\longrightarrow S$, i.e., the following fibre product (fibre product of classical Artin stacks)
[…]
The fibre product is usually non-reduced. Let $\mathcal{N}ilp^{red}$ be the reduction of $\mathcal{N}ilp$. It is a stack over the spectrum of a discrete valuation ring $S$. We denote its closed fibre by $\mathcal{N}ilp^{red}_{0}$.
Remark 6.2.
A point in $\mathcal{N}ilp^{red}_{0}$ is a tuple $(X_{r},\mathcal{E},\phi)$, where $X_{r}$ is a Gieseker curve with $r$ $\mathbb{P}^{1}$ bubbles $(0\leq r\leq n)$, $\mathcal{E}$ is a Gieseker vector bundle on $X_{s}$ and $\phi$ is a nilpotent Higgs field, i.e., $\phi^{n}=0$. Given a nilpotent Higgs field, we get a canonical filtration by saturated torsion-free sub-sheaves
$\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$ (6.2)
for some integer $k$ such that $\phi^{k+1}=0$.
Definition 6.3.
(Type of a nilpotent torsion-free Higgs pair on the nodal curve $X_{0}$) Let $(\mathcal{F},\psi)$ be a torsion-free Higgs pair on $X_{0}$, where $\mathcal{F}$ is a torsion-free coherent sheaf on $X_{0}$ and $\psi:\mathcal{F}\longrightarrow\mathcal{F}\otimes\omega_{X_{0}}$ is a map of coherent sheaves. Suppose that the Higgs field is nilpotent, i.e., $\psi^{n}=0$, where $n$ denotes the rank of the torsion-free sheaf $\mathcal{F}$. Then, as mentioned above, we get a natural flag of saturated sub sheaves of $\mathcal{F}$.
$\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$ (6.3)
For every $11\leq i\leq k+$ 1, we define $n_{i}:=\textsf{rank}\leavevmode\nobreak\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$ and $d_{i}:=\deg\leavevmode\nobreak\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$, where the degree of a torsion-free sheaf on an irreducible projective curve is defined to be its first Chern number. We say that the nilpotent Higgs bundle $(\mathcal{F},\psi)$ is of type $\{(n_{i},d_{i})\}^{i=k+1}_{i=1}$. We denote it by $\textsf{Type}\leavevmode\nobreak\ (\mathcal{F},\psi)$.
Definition 6.4.
(Type of a nilpotent Higgs bundle on a Gieseker curve $X_{r}$ for some $r\in[0,n]$) Let $r\in[0,n]$ be an integer. Let $\pi_{r}:X_{r}\longrightarrow X_{0}$ be the Gieseker curve with exactly $r$ many $\mathbb{P}^{1}$’s. Let $(\mathcal{E},\phi)$ be a Gieseker-Higgs bundle on $X_{r}$. Then, by definition, $((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$ is a nilpotent torsion-free Higgs pair on $X_{0}$. We define $\textsf{Type}\leavevmode\nobreak\ (X_{r},\mathcal{E},\phi):=\textsf{Type}\leavevmode\nobreak\ ((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$.
Lemma 6.5.
Let $(\pi_{r}:X_{r}\rightarrow X_{0},\mathcal{E},\phi)$ be a nilpotent Gieseker-Higgs bundle. Then we have the following induced filtration as in (6.2).
$\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$ (6.4)
The induced nilpotent torsion-free Higgs pairs $(\mathcal{F}:=(\pi_{r})_{*}\mathcal{E},\psi:=(\pi_{r})_{*}\phi)$ also has a natural filtration as in (6.3).
$\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$ (6.5)
Then $(\pi_{r})_{*}\mathcal{E}^{i}\cong\mathcal{F}^{i}$ for every $11\leq i\leq k+$ 1.
Proof.
For every $i\in[1,k+1]$, we have morphisms $\phi^{i}:\mathcal{E}\longrightarrow\mathcal{E}\otimes\omega^{\otimes i}_{X_{r}}$ and $\psi^{i}:\mathcal{F}\longrightarrow\mathcal{F}\otimes\omega^{\otimes i}_{X_{0}}$. Moreover, $\ker(\phi^{i})=\mathcal{E}^{i}$ and $\ker(\psi^{i})=\mathcal{F}^{i}$. Let $\sigma$ be a local section of $(\pi_{r})_{*}\mathcal{E}^{i}$ on a neighbourhood $U$ of the node of $X_{0}$. It is an element of $\mathcal{E}^{i}((\pi_{r})^{-1}(U))$. Therefore the section $\phi^{i}(\sigma)=0$ on the open set $(\pi_{r})^{-1}(U)\cap\tilde{X_{0}}=U^{o}$, where $U^{o}$ denotes the complement of the node of $X_{0}$. Therefore $\psi^{i}(\sigma)=0$ on $U^{o}$, because $\psi=\phi$ on $U^{o}$. Since $U^{o}$ is dense in $U$, therefore $\psi^{i}(\sigma)=0$ on $U$ and $\sigma\in\mathcal{F}^{i}$. Therefore, $(\pi_{r})_{*}\mathcal{E}^{i}\subseteq\mathcal{F}^{i}$ for all $i\in[1,k+1]$.
For the converse, let $\sigma\in\mathcal{F}^{i}(U)$. Since $\mathcal{F}^{i}\subseteq\mathcal{F}=(\pi_{r})_{*}\mathcal{E}$, we have $\sigma\in\mathcal{E}((\pi_{r})^{-1}(U))$. Since $\psi^{i}(\sigma)=0$, therefore $\phi^{i}(\sigma)=0$ on $(\pi_{r})^{-1}(U^{o})$ and hence, by continuity, on $(\pi_{r})^{-1}(U)\cap\widetilde{X_{0}}$. Now notice that $\phi^{i}(\sigma)$ is a section of $\mathcal{E}\otimes\omega_{X_{r}}$. The bundle $\mathcal{E}\otimes\omega_{X_{r}}$ is a Gieseker-vector bundle because $\omega_{X_{r}}|_{R}\cong\mathcal{O}_{R}$, where $R$ the chain of $\mathbb{P}^{1}$’s in $X_{r}$. Therefore, it must vanish everywhere since the section vanishes at the two points $\widetilde{X_{0}}\cap R$. This implies that $\sigma\in\mathcal{E}^{i}((\pi_{r})^{-1}(U))=((\pi_{r})_{*}\mathcal{E}^{i})(U)$. Therefore, $\pi_{r*}(\mathcal{E}^{i})=\mathcal{F}^{i}$ for all $i\in[1,k+1]$. ∎
Definition 6.6.
We define $\mathcal{N}ilp^{sm,gen}_{0}$ to be the open substack of $\mathcal{N}ilp^{red}_{0}$ consisting of nilpotent Higgs bundles $(X_{r},\mathcal{E},\phi)$ which satisfy the following two conditions.
  1. 1.
    $(X_{r},\mathcal{E},\phi)$ is a smooth point of $\mathcal{N}ilp^{red}_{0}$.
  2. 2.
    the $\textsf{Type}(X_{r},\mathcal{E},\phi)$ is the same as the Type of the generic point of the component of the reduced nilpotent cone that $(X_{r},\mathcal{E},\phi)$ belongs to.
Let $(\mathcal{X}^{mod}_{{}_{uni}},\mathcal{E}_{{}_{uni}},\phi_{{}_{uni}})$ denote the restriction of the universal modification, universal vector bundle and the universal Higgs field to $M^{cl}_{Gie}$. We restrict it to $\mathcal{N}ilp^{sm,gen}_{0}$. For simplicity of the notation, let us drop the subscript ”univ” from the notation and simply denote it by $(\mathcal{X}^{mod},\mathcal{E},\phi)$.
Lemma 6.7.
Consider the filtration 6.2 of $\mathcal{E}$ induced by the Higgs field $\phi$. Then the sheaves $\mathcal{E}^{i}$ in the filtration are all flat over $\mathcal{N}ilp^{sm,gen}_{0}$.
Proof.
Consider any connected component $C$ of $\mathcal{N}ilp^{sm,gen}_{0}$ to see this. Then we have the following exact sequence for any element $c:{\rm Spec}\,\mathbb{k}\longrightarrow C$.
$0\longrightarrow\textsf{K}_{c}^{i}\longrightarrow c^{*}\mathcal{E}\xrightarrow{(c^{*}\phi)^{i}}c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathcal{X}^{mod}}\longrightarrow\textsf{CK}^{i}_{c}\longrightarrow 0$
Here $\textsf{K}_{c}^{i}$ and $\textsf{CK}^{i}_{c}$ denote the kernel and cokernel of the map $(c^{*}\phi)^{i}$, respectively. We also have the following exact sequence.
$0\longrightarrow\underline{\textsf{K}}^{i}_{c}\longrightarrow(\pi_{r})_{*}(c^{*}\mathcal{E})\xrightarrow{(\pi_{r})_{*}((c^{*}\phi)^{i})=((\pi_{r})_{*}(c^{*}\phi))^{i}}(\pi_{r})_{*}(c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathcal{X}^{mod}})\longrightarrow\underline{\textsf{CK}}^{i}_{c}\longrightarrow 0$
Here $\underline{\textsf{K}}^{i}_{c}$ and $\underline{\textsf{CK}}^{i}_{c}$ are kernel and cokernel, respectively. Notice from Lemma 6.5 it follows that $\underline{\textsf{K}}^{i}_{c}=(\pi_{s})_{*}(\textsf{K}^{i}_{c})$. Since the $\textsf{Type}\leavevmode\nobreak\ \leavevmode\nobreak\ (c^{*}{\mathcal{X}^{mod}},c^{*}\mathcal{E},c^{*}\phi)$ is constant over $C$, therefore the Hilbert polynomial of $\underline{\textsf{K}}^{i}_{c}$ does not depend on $c\in C$ which in turn implies that the Hilbert polynomial of $\underline{\textsf{CK}}^{i}_{c}$ does not depend on $c\in C$. Consider next the sheaf cokernel $\underline{\textsf{CK}}^{i}:=Coker((\pi_{r})_{*}((\phi)^{i})$. Since $\underline{\textsf{CK}}^{i}$ is a cokernel, we have that $c^{*}\underline{\textsf{CK}}^{i}=\underline{\textsf{CK}}^{i}_{c}$. Therefore the sheaf $\underline{\textsf{CK}}^{i}$ is flat over $C$. Now consider the kernel sheaf $\underline{\textsf{K}}^{i}:=Ker((\pi_{r})_{*}((\phi)^{i}))$. Since in the four terms sequence
$0\longrightarrow\underline{\textsf{K}}^{i}\longrightarrow(\pi_{r})_{*}(\mathcal{E})\xrightarrow{(\pi_{r})_{*}((\phi)^{i})=((\pi_{r})_{*}(\phi))^{i}}(\pi_{r})_{*}(\mathcal{E}\otimes\omega^{\otimes i}_{\mathcal{X}^{mod}})\longrightarrow\underline{\textsf{CK}}^{i}\longrightarrow 0$
the last three terms are flat over $C$, we conclude that $\underline{\textsf{K}}^{i}$ will also be flat over $C$, and therefore $\textsf{K}^{i}$ is flat over $C$. ∎
Proposition 6.8.
The tangent complex of $\mathcal{N}ilp^{red}_{0}$ at a point $(X_{r},E,\phi)\in\mathcal{N}ilp^{sm,gen}_{0}$ is given by $R\Gamma(\mathcal{SC}{(E,\phi)})$, where $\mathcal{SC}{(E,\phi)}$ is the following complex of sheaves on $X_{r}$.
$\mathcal{SC}{(E,\phi)}:=[SC(E,\phi)\xrightarrow{[-,\phi]}SC(E,\phi)\otimes\omega_{X_{r}}]$ (6.6)
where $SC(E,\phi)\subseteq\mathcal{E}ndE$ is the sheaf of local sections $s$ of $\mathcal{E}ndE$ such that $s(E^{i})\subseteq E^{i-1}$ for $i=1,\dots,k+1$.
Proof.
We choose a trivialisation $X_{r}=\cup_{i\in I}V_{i}$ of the vector bundle $\mathcal{E}$ and the line bundle $\omega_{X_{r}}$. Then a first-order infinitesimal deformation (as a Higgs bundle) of $(E,\phi)$ can be described as a pair $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}ndE)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}ndE\otimes\omega_{X_{r}})$ satisfying $s_{ij}+s_{jk}=s_{ik}$ and $t_{i}-t_{j}=[s_{ij},\phi]$ (see [5, Theorem 2.3] and [6, Proposition 3.1.2]).
Let $\operatorname{Spec}(\mathbb{k}[\epsilon])\longrightarrow\mathcal{N}ilp^{sm,gen}_{0}$ be a map such that the image of the closed point is given by the Higgs bundle $(X_{r},E,\phi)$. Let us denote by the corresponding first-order infinitesimal deformation of the nilpotent Higgs bundle $(E,\phi)$ by $(E[\epsilon],\phi[\epsilon])$. We assume that $(E[\epsilon],\phi[\epsilon])$ is a nilpotent Higgs bundle. We define $E^{i}[\epsilon]:=\ker\leavevmode\nobreak\ (\phi[\epsilon])^{i}$ for $i\in[1,k+1]$. Since the induced flag $E^{\bullet}$ is flat over $\mathcal{N}ilp^{sm,gen}_{0}$ (see Lemma 6.7), therefore $E^{\bullet}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$. Since $E^{i}[\epsilon]$ is flat over $\operatorname{Spec}(\mathbb{k}[\epsilon])$, therefore it is an extension of $E^{i}$ by $E^{i}$.
[…]
Now it is straightforward to check that for $(s_{ij},t_{i})$ to be an infinitesimal deformation of $(E,\phi)$ as a nilpotent Higgs field, it has to satisfy the extra condition that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$, where $E^{\bullet}$ is the flag (6.2). This means that $s_{ij}\in\Gamma(V_{ij},SC(E,\phi))$ for all $i,j$ and therefore, the proposition follows. ∎
Theorem 6.9.
  1. 1.
    The Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ is surjective.
  2. 2.
    The sub-stack $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic in $M^{cl}_{Gie}$ with respect to the relative 0-shifted log symplectic form which was defined in Theorem 4.9.
  3. 3.
    The Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ is flat.
Proof.
Proof of (1): To prove the surjectivity we use the spectral correspondence [1, Lemma 2.4]. Given any point $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$, we consider the function $s(a_{\bullet}):=t^{n}+a_{1}t^{n-1}+\dots+a_{n-1}t+a_{n}$ on the total space $\operatorname{Tot}(\omega_{X_{0}})$. Here $t$ denotes the canonical section of $f^{*}\omega_{X_{0}}$, where $f:\operatorname{Tot}(\omega_{X_{0}})\longrightarrow X_{0}$ is the projection map. Then the vanishing locus $V(s(a_{\bullet}))$ defines a closed sub-scheme of $\operatorname{Tot}(\omega_{X_{0}})$. Notice that $\operatorname{Tot}(\omega_{X_{0}})$ is only quasi-projective and it is an open subscheme of $Z:=\mathbb{P}(\omega_{X_{0}}^{*}\oplus\mathcal{O}_{X_{0}})$. But since $s(a_{\bullet})$ is a monic polynomial, therefore the closure of $V(s(a_{\bullet}))$ in $Z$ is $V(s(a_{\bullet}))$ itself. In particular, $V(s(a_{\bullet}))$ is a closed subscheme in $Z$ such that $V(s(a_{\bullet}))\cap D_{\infty}=\emptyset$, where $D_{\infty}:=Z\setminus\operatorname{Tot}(\omega_{X_{0}})$. Therefore, by spectral correspondence, it follows that any rank $1$ locally free sheaf on $V(s(a_{\bullet}))$ corresponds to a Higgs bundle $(E,\phi)$ on the nodal curve $X_{0}$ whose characteristic polynomial is given by $a_{\bullet}:=(a_{1},\dots,a_{n})\in B$. Therefore, the Hitchin map is surjective.
Proof of (2): We want to show that $\mathcal{N}ilp^{sm,gen}$ is relatively isotropic over $S$. First, we notice that the smooth locus of the generic fibre of $\mathcal{N}ilp^{red}\longrightarrow S$ is isotropic, and its dimension is equal to $n^{2}(g-1)$, which is the same as the dimension of the stack of rank $n$ vector bundles on a smooth projective curve of genus $g$. Therefore, we have $\dim\mathcal{N}ilp^{red}_{0}\geq n^{2}(g-1)$. We want to show that the $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack in $M^{cl}_{Gie,0}$. The tangent complex at a point $(X_{r},E,\phi)$ of $\mathcal{N}ilp^{red,sm}_{0}$ is given by the derived global sections of (6.6). The tangent complex at the corresponding point of $M^{cl}_{Gie}$ is given by the derived global sections of the following complex (see 4.17 and Theorem 4.18).
$\mathcal{C}{(E,\phi)}:=[\mathcal{E}ndE\xrightarrow{[-,\phi]}\mathcal{E}ndE\otimes\omega_{X_{r}}]$ (6.8)
We want to show that the following composite morphism is homotopic to 00.
${R\Gamma(\mathcal{SC}(E,\phi))[1]}$${R\Gamma(\mathcal{C}(E,\phi))[1]}$${(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}}$${(R\Gamma(\mathcal{SC}(E,\phi))[1])^{\vee}}$$\scriptstyle{\omega^{\flat}}$ (6.9)
Notice that the morphism $\omega^{\flat}:R\Gamma(\mathcal{C}(E,\phi))[1]\to(R\Gamma(\mathcal{C}(E,\phi))[1])^{\vee}$ is given by the log-symplectic pairing (Theorem 4.9). If we show that the composite morphism 6.9 is homotopic to 00, this will imply that the $\mathcal{N}ilp^{red,sm}_{0}$ is an isotropic substack. To show this, we choose a cover $X_{r}=\cup_{i\in I}V_{i}$ for which the vector bundle $E$ and the line bundle $\omega_{X_{r}}$ trivialize. With respect to this cover, the complex $R\Gamma(\mathcal{SC}(E,\phi))[1]$ is quasi-isomorphic to the following Čech complex
$\prod SC(E,\phi)(V_{i})\longrightarrow\prod SC(E,\phi)(V_{ij})\oplus\prod(SC(E,\phi)\otimes\omega_{X_{r}})({V_{i}})\longrightarrow\prod(SC(E,\phi)\otimes\omega_{X_{r}})(V_{ij})\oplus\prod SC(E,\phi)(V_{ijk})$,
where the term $\prod SC(E,\phi)({V_{i}})$ is in degree $-1$.
Notice that the composite map can only be non-zero on three cohomologies: in the degree $-1$, 00 and $1$.
Case 1: the induced map $H^{0}(\omega^{\flat})$: An element of
$\prod SC(E,\phi)(V_{ij})\oplus\prod(SC(E,\phi)\otimes\omega_{X_{r}})({V_{i}})$
is given by $\{(s_{ij},t_{i})\}_{i,j\in I}$, where $(s_{ij},t_{i})$, where $s_{ij}\in\Gamma(V_{ij},\mathcal{E}ndE)$ and $t_{i}\in\Gamma(V_{i},\mathcal{E}ndE\otimes\omega_{X_{r}})$ such that $s_{ij}(E^{\bullet})\subseteq E^{\bullet-1}$, where $E^{\bullet}$ is the flag (6.2). Then from (6.9) it follows that the pairing of two elements $\{(s_{ij},t_{i})\}$ and $\{(s^{\prime}_{ij},t^{\prime}_{i})\}$ is given by
$\{T_{ij}:=Trace(s_{ij}\circ t^{\prime}_{j}-t_{i}\circ s^{\prime}_{ij})\in\omega_{X_{r}}(V_{ij})\}_{i,j\in I}.$
Claim: The section $T_{ij}$ vanishes for all $i,j\in I$.
Proof of the claim: We first observe that, we can choose the trivialisation $X_{r}=\cup_{i\in I}V_{i}$ of $E$ and $\omega_{X_{r}}$ in such a way that
  1. 1.
    $V_{i}$ contains at most one node for each $i\in I$,
  2. 2.
    $V_{i}$’s are connected.
We notice that $V_{ij}$’s also contain at most one node. Suppose, $V_{ij}$ contains a node $p$. Then $V_{ij}=V^{1}_{ij}\coprod V^{2}_{ij}$, the union of two smooth irreducible components of $V_{ij}$. We denote by $V^{o}_{ij}:=V_{ij}\setminus p$, $V^{1,o}_{ij}:=V^{1}_{ij}\setminus p$ and $V^{2,o}_{ij}:=V^{2}_{ij}\setminus p$. We consider the restriction of the section $T_{ij}$ to the two open subsets $V^{1,o}_{ij}$ and $V^{2,o}_{ij}$. Notice that the restrictions of the flag (6.2) to these two open subsets are all sub-bundles. We notice that $SC(E|_{V^{1,o}_{ij}},\phi)$ is the nilpotent part of the parabolic sub-algebra of $\mathcal{E}nd(E|_{V^{1,o}_{ij}})$ given by the natural flag of sub-bundles induced by $\phi$. Therefore it is clear that the trace pairing is 00 i.e., $(T_{ij})|_{V^{1,o}_{ij}}=0$. Similarly, $(T_{ij})|_{V^{2,o}_{ij}}=0$. Since $V^{o}_{ij}$ is a dense open subset of $V_{ij}$, therefore, by continuity, $T_{ij}=0$.
If a $V_{ij}$ does not contain any node, then the proof is similar.
Case 2: the induced map $H^{-1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{i}\in SC(E,\phi)\otimes\omega_{X_{r}}\}$ and $\{s_{ij}\in SC(E,\phi)\}$. The pairing is given by $\{\operatorname{tr}(t_{i}\circ s_{ij})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{i}\circ s_{ij})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i,j\in I$.
Case 3: the induced map $H^{1}(\omega^{\flat})$: In this case, we look at the pairing of $\{t_{ij}\in(SC(E,\phi)\otimes\omega_{X_{r}})(V_{ij})\}$ and $\{s_{j}\in SC(E,\phi)(V_{j})\}$. The pairing is given by $\{\operatorname{tr}(t_{ij}\circ s_{j})\}$. By similar arguments as above, we can show that $\operatorname{tr}(t_{ij}\circ s_{j})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i,j\in I$.
Therefore, we conclude that the smooth locus of this particular component is isotropic.
Proof of (3) From (2), it follows that $\dim\mathcal{N}ilp^{red}_{0}=n^{2}(g-1)$. Now from Lemma 8.11, it follows that the dimension of $M^{cl}_{Gie}$ is equal to $22n^{2}(g-1)+$ 2. Once we know the dimension, it is easy to see that $M^{cl}_{Gie}$ is a local complete intersection (Remark 8.13). Since $M^{cl}_{Gie}$ is a local complete intersection, and all the fibres of the Hitchin map $h:M^{cl}_{Gie}\longrightarrow B$ are of dimension equal to $n^{2}(g-1)$, therefore by the miraculous flatness criterion [13, Exercise 10.9, page 276], it follows that the Hitchin map is flat. ∎
7 On the relative logarithmic Dolbeault moduli over $\overline{\mathcal{M}_{g}}$
In this section, we will construct the derived moduli stack $M_{g}^{Dol}$ of Higgs bundles with coefficients in the relative dualizing sheaf along the fibers of the universal curve over the moduli stack $\mathcal{M}_{g}^{ss}$ of semistable curves of genus $g\geq 2$. We will show that there is a relative 00-shifted symplectic form on $M_{g}^{Dol}/\mathcal{M}_{g}^{ss}$ which can also be interpreted as a relative 00-shifted log-symplectic form on $M_{g}^{Dol}$, relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. Let us first recall some notations.
A connected projective variety $C$ of dim $1$ is called a stable curve if it is either smooth or has nodal singularities, and the automorphism group $\operatorname{Aut}(C)$ is finite.
The moduli stack of stable curves represents the moduli problem
$\overline{\mathcal{M}_{g}}:(Sch/\mathbb{k})^{op}\longrightarrow Groupoids$
$T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of stable curves}\\ \text{ of genus}\leavevmode\nobreak\ g\leavevmode\nobreak\ \text{over}\leavevmode\nobreak\ T\end{array}\right\}$ (7.1)
Here, morphisms in the groupoid $\overline{\mathcal{M}_{g}}(T)$ are isomorphisms of stable curves over $T$.
Remark 7.1.
It is well known that if $g\geq 2$, then $\overline{\mathcal{M}_{g}}$ is a smooth and proper Deligne-Mumford stack. The locus of nodal curves forms a normal crossing divisor in $\overline{\mathcal{M}_{g}}$ and so gives a divisorial log structure on $\overline{\mathcal{M}_{g}}$. On the other hand, by a well known result of S. Mochizuki [23, §3B.], given any family of semi-stable curves $\mathcal{C}\to T$, there is a natural log structure on $\mathcal{C}$ and also on the base scheme $T$ such that the projection map is a log-smooth map and the relative logarithmic cotangent bundle is the relative dualizing sheaf of $\mathcal{C}/T$. We call these log structures the basic log structure package on $\mathcal{C}/T$. It is known that the log structure on $\overline{\mathcal{M}_{g}}$ which is part of the basic log structure package for the universal curve $\mathcal{C}_{g}$ over $\overline{\mathcal{M}_{g}}$ agrees with the divisorial log-structure on $\overline{\mathcal{M}_{g}}$ (see the comment after Lemma 4.4 in [21] and the proof of our Proposition 7.3).
A connected projective variety $C$ of dim $1$ is called a semi-stable curve if the following properties are satisfied
  1. 1.
    it is either smooth or has nodal singularities,
  2. 2.
    every rational irreducible component $R$ is smooth, and
  3. 3.
    $R\cdot\overline{(C\setminus R)}=2$.
The moduli stack of semi-stable curves represents the moduli problem
${\mathcal{M}^{ss}_{g}}:Sch/\mathbb{k}\longrightarrow Groupoids$
$T\mapsto\left\{\begin{array}[]{@{}ll@{}}\text{families of semi-stable curves}\\ \text{of genus}\leavevmode\nobreak\ g\leavevmode\nobreak\ \text{over}\leavevmode\nobreak\ T\end{array}\right\}$ (7.2)
Morphisms in the groupoid ${\mathcal{M}^{ss}_{g}}(T)$ are again isomorphisms of families of semi-stable curves over $T$.
Remark 7.2.
It is well known that if $g\geq 2$, ${\mathcal{M}^{ss}_{g}}$ is a smooth Artin stack [39, 0E72, Lemma 21.5] of finite type.
7.1 Log structures on ${\mathcal{M}^{ss}_{g}}$
The locus of singular curves in ${\mathcal{M}^{ss}_{g}}$ again forms a normal crossing divisor, and hence induces a log structure on ${\mathcal{M}^{ss}_{g}}$. We denote this divisor by $\partial\mathcal{M}^{ss}_{g}$. Let us denote the universal curve over ${\mathcal{M}^{ss}_{g}}$ by $\mathcal{D}_{g}$. There is the “stabilization morphism” $\pi:\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ such that the induced morphism $\mathcal{D}_{g}\longrightarrow\pi^{*}\mathcal{C}_{g}$ is the universal modification morphism over $\mathcal{M}^{ss}_{g}$.
7.1.1 Versal deformations of the stabalization morphism
Let $\mathcal{X}$ be a stable curve of genus $g$ over ${\rm Spec}\,\mathbb{k}$ and $\mathfrak{X}$ be a semi-stable curve whose stable model is $\mathcal{X}$. Let $\{c_{i}\}^{l}_{i=1}$ be the nodes of the curve $\mathcal{X}$. Let $\{d_{ij}\}^{\iota_{i}}_{i=1}$ be the nodes of $\mathfrak{X}$ over the node $c_{i}$ for every $i=1,\dots,l$. From [35, Proposition 3.3.2], it follows that
  1. 1.
    there exists a versal deformation space of the nodal curve $\mathcal{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|z_{1},\dots,z_{l}|][|z_{l+1},\dots,z_{N}|]$, where $N:=3g-3$ (here $g$ is the genus of the curves) and $z_{i}$ is the equation of the $i$-th node of $\mathcal{X}$ for $i=1,\dots,l$, and
  2. 2.
    there exits a versal deformation space of the nodal curve $\mathfrak{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}|][|z_{l+1},\dots,z_{N}|]$, where $z_{ij}=0$ is the equation of the node $d_{ij}$ for $i=1,\dots,l$ and $j=1,\dots,\iota_{i}$. Moreover, there is a morphism between the versal deformation spaces, which given is as follows.
    $\mathbb{k}[|z_{1},\dots,z_{l}|][|z_{l+1},\dots,z_{N}|]\longrightarrow\mathbb{k}[|\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}|][|z_{l+1},\dots,z_{N}|]$
    $z_{i}\mapsto z_{i1}\cdots z_{i\iota_{i}}\hskip 30.00005pt\forall i=1,\dots,l,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}$
    $z_{j}\mapsto z_{j}\hskip 30.00005pt\forall j=l+1,\dots,N$
Proposition 7.3.
  1. 1.
    The log structures on $\overline{\mathcal{M}_{g}}$ and $\mathcal{M}^{ss}_{g}$ which are parts of the basic log structure package on $\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}$ and $\mathcal{D}_{g}/\mathcal{M}^{ss}_{g}$ (see Remark 7.1) are locally free. These log structures coincide with the divisorial log structures induced by the boundary divisors $\partial\overline{\mathcal{M}_{g}}$ and $\partial\mathcal{M}^{ss}_{g}$, respectively.
  2. 2.
    The morphism $\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ is a log-smooth morphism with respect to the log-structures mentioned in (1).
Proof.
Proof of (1): The fact that the log structure from the basic log structure packages and the log structures induced from the boundary divisors are the same can be seen as follows.
Given a point $p\in\overline{\mathcal{M}_{g}}$ and the set of nodes $\{q_{1},\dots,q_{n}\}\in\mathcal{C}_{g,p}$ (the fiber over $p$ of the universal curve $\mathcal{C}_{g}$), we have a $\mathbb{k}$ algebra $B$ and an element $b_{i}$ fitting into a commutative diagram of étale neighbourhoods of the points $p$ and $q_{i}\in\{q_{1},\dots,q_{n}\}$ (using Artin approximation):
[…]
Here the rightmost square is cartesian and $U_{i}$ denotes a suitable étale neighbourhood of $q_{i}$ and not containing any other nodes in $\{q_{1},\dots,q_{n}\}$. The scheme ${\rm Spec}\,\mathbb{k}[a]$ is a versal deformation space of a node (i.e., ${\rm Spec}\,\mathbb{k}[x,y]/(xy$)) and the family
${\rm Spec}\,(\mathbb{k}[x,y,a]/(xy-a))\to{\rm Spec}\,\mathbb{k}[a]$
is a versal deformation family of curves for a node (see [14, Theorem 14.1]). If we consider the basic log structure package on ${\rm Spec}\,(\mathbb{k}[x,y,a]/(xy-a))/\operatorname{Spec}\leavevmode\nobreak\ (\mathbb{k}[a])$, the basic log structures on $\operatorname{Spec}\leavevmode\nobreak\ \mathbb{k}[a]$, agrees by Kato [21, Lemma 4.4] with the divisorial log structure given by the boundary divisor (i.e., the vanishing locus of the element $a\in\mathbb{k}[a]$). Now from [21, Lemma 2.1, Lemma 2.2, Proposition 2.3], we can find an étale neighbourhood $V$ of the point $p\in\overline{\mathcal{M}_{g}}$, and define a log structure by taking the amalgamated sum of the log structures contributed by each of the nodes in $\{q_{1},\dots,q_{n}\}$. But this is the same as the log structures induced by the pullback of the boundary divisor in $\overline{\mathcal{M}_{g}}$. This completes the argument. Similarly, the two log structures on $\mathcal{M}^{ss}_{g}$ are the same.
Proof of (2): Again using Artin approximation we have a commutative diagram
[…]
where all the squares and triangles are commutative and only the left square is Cartesian.
It can be rewritten as the following commutative diagram
[…]
Notice that the rightmost vertical morphism is a morphism of log schemes. Since the right commutative diagram is commutative and the log structures on $V$ and $W$ are isomorphic to the pullback of the log structures from $\operatorname{Spec}\leavevmode\nobreak\ \mathbb{k}[\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}][\{z_{k}\}^{N}_{k=l+1}]$ and $\mathbb{k}[z_{1},\dots,z_{l}][z_{l+1},\dots,z_{N}]$, therefore we have the structure of a log-morphism on $V\rightarrow W$.
Now we cover $\mathcal{M}^{ss}_{g}$ and $\overline{\mathcal{M}_{g}}$ with compatible étale open subsets with log morphisms between them. Now we need to show that these log-morphisms glue i.e., they are equal on the intersections. But this can be checked by passing to the completion, where the log structure is given by the following chart:
[…]
where $e_{i}$ is sent to $z_{i}$ and $e_{ij}$ is sent to $z_{ij}$, where $i=1,\dots,l$. ∎
7.2 Relative log-cotangent complex of the map ${\mathcal{M}^{ss}_{g}}\stackrel{{\scriptstyle f}}{{\longrightarrow}}\overline{\mathcal{M}_{g}}$
Proposition 7.4.
The relative logarithmic cotangent complex of $f$ is trivial: $\mathbb{L}^{\log}_{f}\cong 0$.
Proof.
We have a commutative diagram (not a cartesian square)
${\mathcal{D}_{g}}$${\mathcal{C}_{g}}$${\mathcal{M}^{ss}_{g}}$${\overline{\mathcal{M}_{g}}}$$\scriptstyle{\tilde{p}}$$\scriptstyle{\tilde{f}}$$\scriptstyle{p}$$\scriptstyle{f}$ (7.7)
These families of semi-stable curves as discussed in Remark 7.1 have relative log structure packages, defining log structures on all the spaces in this diagram, for which the projection morphisms $p$ and $\tilde{p}$ are log-smooth. The log-tangent complex of $\overline{\mathcal{M}_{g}}$ is isomorphic to $Rp_{*}(\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}})[1]$ and the log-tangent complex of $\mathcal{M}^{ss}_{g}$ is isomorphic to $R\tilde{p}_{*}(\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}})[1]$. Now notice that $\tilde{f}^{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$ and that the fibers of $\tilde{f}$ are connected and so $R\tilde{f}_{*}\mathcal{O}\cong\mathcal{O}$. Hence, $R\tilde{f}_{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$. Therefore we have,
$Rf_{*}\circ R\tilde{p}_{*}(\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}})[1]=Rp_{*}R\tilde{f}_{*}(\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}})[1]\cong Rp_{*}(\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}})[1].$
Therefore, we conclude that the natural map between the two log-tangent complexes ( and the map between log-cotangent complexes) is an isomorphism. We have the following distinguished triangles of log-cotangent complexes
$f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{{\mathcal{M}^{ss}_{g}}}\longrightarrow\mathbb{L}^{\log}_{f}\longrightarrow f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}[1]$ (7.8)
of the map
$f:\mathcal{M}^{ss}_{g}\to\overline{\mathcal{M}_{g}}.$
Since, the map $f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{{\mathcal{M}^{ss}_{g}}}$ is an equivalence, therefore we conclude that $\mathbb{L}^{\log}_{f}\cong 0$. ∎
7.3 Relative logarithmic Dolbeault shape and shifted symplectic forms
Let $\mathcal{D}^{Dol}_{g}$ denote the relative logarithmic Dolbeault moduli stack for the family of curves $\mathcal{D}_{g}\longrightarrow\mathcal{M}^{ss}_{g}$. Then $\mathcal{M}^{Dol}_{g}:=\mathsf{Map}_{\mathcal{M}^{ss}_{g}}(\mathcal{D}^{Dol}_{g},BGL_{n}\times\mathcal{M}^{ss}_{g})$ is the relative derived moduli stack of Gieseker-Higgs bundles viewed over $\overline{\mathcal{M}_{g}}$.
Proposition 7.5.
The morphism $\mathcal{M}^{Dol}_{g}\longrightarrow\mathcal{M}^{ss}_{g}$ is a quasi-smooth morphism of derived Artin stacks.
Proof.
Same as the proof of Proposition 4.8. ∎
We equip $\mathcal{M}^{Dol}_{g}$ with the locally free log-structure pulled back from $\mathcal{M}^{ss}_{g}$ via the morphism $\mathcal{M}^{Dol}_{g}\longrightarrow\mathcal{M}^{ss}_{g}$.
Theorem 7.6.
There is a 00-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
Proof.
[…]
$\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}[1].$ (7.9)
But since $\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\cong 0$ (Proposition 7.4), therefore the relative log cotangent complex of the composite morphism is isomorphic to the relative log cotangent complex of the morphism $\pi$. But notice that the log structure of $\mathcal{M}^{Dol}_{g}$ is pulled back of the log structure of $\mathcal{M}^{ss}_{g}$ via the map $\pi$. Therefore the relative log-cotangent complex of the morphism $\pi$ is isomorphic to the relative cotangent complex of the morphism $\pi$. Now since the relative logarithmic Dolbeault stack over the moduli stacks $\mathcal{M}^{Dol}_{g}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented (Theorem 4.7); therefore $\mathcal{M}^{Dol}_{g}$ has a 00-shifted relative symplectic form over $\mathcal{M}^{ss}_{g}$, which is a 00-shifted relative log-symplectic form viewed over the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. ∎
8 Appendix: Classical Artin stack of Gieseker-Higgs bundles and its local properties
In this section, we will define and construct the relative classical Artin stack of Gieseker-Higgs bundles and study its local properties. The main results of the appendix are Lemma 8.10 and Lemma 8.11. In the first lemma, we prove that the stack of Gieseker vector bundles is an almost very good stack ([37, Definition 2.1.2]). We use this in Lemma 8.11 to show that the classical stack of Gieseker-Higgs bundles is an irreducible, local complete intersection.
8.1 The stacks of Gieseker vector bundles and Gieseker Higgs bundles
Definition 8.1.
(Gieseker vector bundle) A vector bundle $\mathcal{E}$ of rank $n$ on $X_{r}$ with $r\geq 1$ is called a Gieseker vector bundle if
  1. 1.
    $\mathcal{E}|_{R[r]}$ is a strictly standard vector bundle on $R[r]\subset X_{r}$, i.e., for each $i=1,\dots,r$, $\exists$ non-negative integers $a_{i}$ and $b_{i}$ such that $\mathcal{E}|_{R[r]_{i}}\cong\mathcal{O}^{\oplus a_{i}}\oplus\mathcal{O}(1)^{\oplus b_{i}}$, and
  2. 2.
    the direct image $\pi_{r*}(\mathcal{E})$ is a torsion-free $\mathcal{O}_{X_{0}}$-module.
Any vector bundle on $X_{0}$ is called a Gieseker vector bundle. In the literature, a Gieseker vector bundle is also called an admissible vector bundle.
A Gieseker vector bundle $(X_{r},\mathcal{E})$ is called a stable Gieseker vector bundle if $\pi_{r*}\mathcal{E}$ is a stable torsion-free sheaf on the irreducible nodal curve $X_{0}$, where $\pi_{r}:X_{r}\longrightarrow X_{0}$ is the natural contraction map.
A (stable) Gieseker vector bundle on a modification $\mathfrak{X}_{T}$ (see Definition 3.2) is a vector bundle such that its restriction to each $(\mathfrak{X}_{T})_{t}$ is a (stable) Gieseker vector bundle.
Remark 8.2.
A Gieseker vector bundle can also be defined as a vector bundle $(X_{r},\mathcal{E})$ on a Gieseker curve $X_{r}$ satisfying the following two conditions
  1. 1.
    $\mathcal{E}|_{R[r]}$ is a globally generated vector bundle on $R[r]$,
  2. 2.
    the direct image $\pi_{r*}(\mathcal{E})$ is a torsion-free $\mathcal{O}_{X_{0}}$-module.
Definition 8.3.
(Gieseker-Higgs bundle) A Gieseker-Higgs bundle on $\mathfrak{X}_{T}$ is a pair $(\mathcal{E}_{T},\phi_{T})$, where $\mathcal{E}_{T}$ is a vector bundle on $\mathfrak{X}_{T}$, and $\phi_{T}:\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}\otimes\omega_{\mathfrak{X}_{T}/T}$ is an $\mathcal{O}_{\mathfrak{X}_{T}}$module homomorphism satisfying the following
  1. 1.
    $\mathcal{E}_{T}$ is a Gieseker vector bundle on $\mathfrak{X}_{T}$,
  2. 2.
    for each closed point $t\in T$ over $\eta_{0}\in S$, the direct image $(\pi_{t})_{*}(\mathcal{E}_{t})$ is a torsion-free sheaf on $X_{0}$ and $(\pi_{t})_{*}\phi_{t}:(\pi_{t})_{*}(\mathcal{E}_{t})\longrightarrow(\pi_{t})_{*}(\mathcal{E}_{t})\otimes\omega_{X_{0}}$ is an $\mathcal{O}_{X_{0}}$-module homomorphism. We refer to such a pair $((\pi_{t})_{*}(\mathcal{E}_{t}),(\pi_{t})_{*}\phi_{t})$ as a torsion-free Higgs pair on the nodal curve $X_{0}$.
Remark 8.4.
A Gieseker-Higgs bundle can also be defined as a Higgs bundle $(X_{r},\mathcal{E},\phi)$ on a Gieseker curve $X_{r}$ satisfying the following two conditions
  1. 1.
    $\mathcal{E}|_{R[r]}$ is a globally generated vector bundle on $R[r]$,
  2. 2.
    the direct image $\pi_{r*}(\mathcal{E})$ is a torsion-free $\mathcal{O}_{X_{0}}$-module.
Remark 8.5.
From [25, Definition-Notation 1, Lemma 2, and Proposition 5], it follows that for the moduli problem of vector bundles (Higgs bundles) of rank $n$, we have to consider Gieseker curves $X_{r}$, where $r=0,\dots,n$.
We fix the rank and degree to be $n$ and $d$, respectively. Also, $n\geq 1$ and $d\in\mathbb{Z}$.
8.1.1 Stack of torsion-free Hitchin pairs
We recall the definition of the moduli stack of torsion-free Hitchin pairs
$TFH(\mathcal{X}/S):{Sch}/S\rightarrow{Groupoids}$
[…]
It is an Artin stack. For the construction of an atlas for $TFH(\mathcal{X}/S)$, we refer to [1, 5.0.7. The total family construction.]. Following the notation from [1], we denote it by $\coprod_{m\geq m_{0}}R^{\Lambda,m}_{S}$. The superscript “$\Lambda$” is because for the construction of the quot scheme of torsion-free Hitchin pairs one views them as $\Lambda$-modules, where $\Lambda=\operatorname{Sym}\leavevmode\nobreak\ (\omega^{\vee}_{\mathcal{X}/S})$ [36]. We denote the stack of families of torsion-free sheaves over $\mathcal{X}/S$ of rank $n$ and degree $d$ by $TF(\mathcal{X}/S)$. The stack $TF(\mathcal{X}/S)$ is a reduced and irreducible Artin stack.
8.1.2 Ordinary Artin Stack of Gieseker-Higgs bundles
Given any derived Artin stack $F:\textsf{cdga}\longrightarrow\mathbb{S}$ one can define a classical/ordinary Artin stack by considering the composition functor $F^{cl}:\textsf{alg}_{\mathbb{k}}\longrightarrow\textsf{cdga}_{\mathbb{k}}\longrightarrow\mathbb{S}$. We call $F^{cl}$ as the underlying classical Artin stack of $F$. Following this notation, we denote the underlying classical Artin stack of the derived stack $M$ of Higgs bundles over the family of curves $\mathcal{X}_{\mathfrak{M}}/\mathfrak{M}$ by $M^{cl}$.
Let us denote by $\operatorname{Coh}(\mathcal{X}/S)$ the Artin stack of coherent $\mathcal{O}_{\mathcal{X}}-$ modules which are flat over $S$. There is a natural map
$\theta:M^{cl}\longrightarrow\operatorname{Coh}(\mathcal{X}/S)$
which is given by the push forward of the underlying bundle $\pi_{*}\mathcal{E}$, where $\pi:\mathcal{X}^{mod}\longrightarrow\mathcal{X}$ is the modification morphism.
Consider the open sub-stack $TF(\mathcal{X}/S)\subset\operatorname{Coh}(\mathcal{X}/S)$ consisting of torsion-free sheaves. The stack $\theta^{-1}(TF(\mathcal{X}/S))$ is an open substack of $M^{cl}$. Let us consider the natural map $\pi^{*}\pi_{*}\mathcal{E}\longrightarrow\mathcal{E}$ over the universal curve $\mathfrak{X}$ over $M^{cl}$. Consider the sub-stack of $M^{cl}$ where the map is surjective. It is again an open sub-stack. Let us denote it by $M^{gg}$ (the superscript ”gg” stands for globally generated). Then we define an open sub-stack
$M^{cl}_{Gie}:=\theta^{-1}(TF(\mathcal{X}/S))\cap M^{gg}$
It is an open sub-stack consisting of Gieseker-Higgs bundles of rank $n$ and degree $d$. There is a natural map
$M^{cl}_{Gie}\longrightarrow TFH(\mathcal{X}/S)$ (8.2)
$(\mathfrak{X},\mathcal{E},\phi)\mapsto(\pi_{*}\mathcal{E},\pi_{*}\phi)$ (8.3)
where $\pi:\mathfrak{X}\longrightarrow\mathcal{X}$ is the modification map. We denote by $M^{{}^{cl}}_{Gie,0}$ the closed fibre of the map $M^{{}^{cl}}_{Gie}\longrightarrow S$.
8.1.3 Construction of an atlas for $N_{Gie}$
Now, we recall the construction of an atlas for the stack $M^{{}^{cl}}_{Gie}$ as well as the classical stack of Gieseker vector bundles. Let us denote by $N_{Gie}$ the classical stack of Gieseker vector bundles.
Let $\mathcal{O}_{\mathcal{X}/S}(1)$ be a relatively ample line bundle for the family of curves $\mathcal{X}/S$. The set of all flat families of stable torsion-free sheaves (Higgs pairs) of degree $d$ and rank $n$ over $\mathcal{X}$ forms a bounded family. Therefore we can choose a large integer $m_{0}$ such that given any family of stable torsion-free Higgs pairs $(\mathcal{F}_{S},\phi_{S})$, the sheaf $\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0})$ is generated by global sections and $H^{1}(\mathcal{X}_{s},\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0}))=0$ for every geometric point $s\in S$. Set $N(m_{0}):=H^{0}(\mathcal{X}_{s},\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0}))$ for any geometric point $s\in S$. We denote by $Grass(N(m_{0}),n)$ the Grassmannian of $n$ dimensional quotient vector spaces of $\mathbb{k}^{N(m_{0})}$.
Definition 8.6.
Let $\mathcal{G}^{m_{0}}_{S}:{Sch}/S\longrightarrow{Sets}$ be the functor defined as follows:
$\mathcal{G}^{m_{0}}_{S}(T)=\{(\Delta_{T},V_{T})\},$ (8.4)
where
$\Delta_{T}\subset\mathcal{X}\times_{S}T\times{Grass}(N(m_{0}),n)$ (8.5)
is a closed subscheme, and $V_{T}$ is a vector bundle on $\Delta_{T}$ such that
  1. 1.
    the projection $j:\Delta_{T}\longrightarrow T\times{Grass}(N(m_{0}),n)$ is a closed immersion,
  2. 2.
    the projection $\Delta_{T}\longrightarrow\mathcal{X}\times_{S}T$ is a modification,
  3. 3.
    the projection $p_{T}:\Delta_{T}\longrightarrow T$ is a flat family of Gieseker curves,
  4. 4.
    Let $\mathcal{V}$ be the tautological quotient bundle of rank $n$ on ${Grass}(N(m_{0}),n)$ and $\mathcal{V}_{T}$ its pullback to $T\times{Grass}(N(m_{0}),n)$. Then
    $V_{T}:=j^{*}(\mathcal{V}_{T})$ (8.6)
    be such that $V_{T}$ is a Gieseker vector bundle on the modification $\Delta_{T}$ of rank $n$ and degree $d(m_{0})^{\prime}:=N(m_{0})+n(g-1)$.
  5. 5.
    for each $t\in T$, the quotient $\mathcal{O}_{\Delta_{t}}^{N}(m_{0})\longrightarrow V_{t}$ induces an isomorphism
    $H^{0}(\Delta_{t},\mathcal{O}_{\Delta_{t}}^{N(m_{0})})\cong H^{0}(\Delta_{t},V_{t})$ (8.7)
    and $H^{1}(\Delta_{t},V_{t})=0$.
We denote by $P(m_{0})$ the Hilbert polynomial of the closed subscheme $\Delta_{s}$ of $\mathcal{X}_{s}\times Grass(N(m_{0}),n)$ for any geometric point $s\in S$ with respect to the polarisation $\mathcal{O}_{\mathcal{X}_{s}}(1)\boxtimes\mathcal{O}_{Grass(N(m_{0}),n)}(1)$, where $\mathcal{O}_{Grass(N(m_{0}),n)}(1)$ is the line bundle $\det\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{V}$.
It is shown in [25, Proposition 8] that the functor $\mathcal{G}^{m_{0}}_{S}$ is represented by a $PGL(N(m_{0}))$-invariant open subscheme $\mathcal{Y}^{m_{0}}_{S}$ of the Hilbert scheme $\mathcal{H}_{S}:={Hilb}^{P(m_{0})}(\mathcal{X}\times{Grass}(N(m_{0}),n))$.
Finally, the disjoint union $\mathcal{Y}_{S}:=(\underset{N(m),m\geq m_{0}}{\coprod}\mathcal{Y}^{m}_{S})\longrightarrow N_{Gie}$ is an atlas. The varieties $\mathcal{Y}^{m}_{S}$ are all smooth varieties and the projection map to the stack is also smooth.
8.1.4 Construction of an atlas for $M^{{}^{cl}}_{Gie}$
Definition 8.7.
We define a functor
$\mathcal{G}^{H,m}_{S}:{Sch}/\mathcal{Y}_{S}\longrightarrow{Groups}$ (8.8)
which maps
$T\longrightarrow H^{0}(T,(p_{T})_{*}(\mathcal{E}nd\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{V}_{T}\otimes\omega_{\Delta_{T}/T})),$
where $p_{T}:\Delta_{T}:=\Delta_{{}_{\mathcal{Y}_{S}}}\times_{{}_{\mathcal{Y}_{S}}}T\longrightarrow T$ is the projection, and $\omega_{\Delta_{T}/T}$ denotes the relative dualising sheaf of the family of curves $p_{T}$.
Since $\mathcal{Y}^{m}_{S}$ is a reduced scheme, the functor $\mathcal{G}^{H,m}_{S}$ is representable, i.e., there exists a linear $\mathcal{Y}^{m}_{S}$- scheme $\mathcal{Y}^{H,m}_{S}$, which represents it. For a $S$-scheme $T$, a point in $\mathcal{G}^{H,m}_{S}(T)$ is given by $(V_{T},\phi_{T})$, where
  1. 1.
    $V_{T}\in\mathcal{G}^{m}_{S}(T)$, and
  2. 2.
    $(V_{T},\phi_{T})$ is a Gieseker-Higgs bundle.
Finally, the disjoint union $\mathcal{Y}^{H}_{S}:=(\underset{N(m),m\geq m_{0}}{\coprod}\mathcal{Y}^{H,m}_{S})\longrightarrow M^{{}^{cl}}_{Gie}$ is an atlas. The projection map to the stack is smooth.
Remark 8.8.
Notice that there is a morphism $\tilde{\theta}:\mathcal{Y}^{H,m}_{S}\to R^{\Lambda,m}_{S}$ given by pushforward
$\tilde{\theta}:\mathcal{Y}^{H,m}_{S}\to R^{\Lambda,m}_{S}$ (8.9)
$(\Delta_{T},V_{T},\phi_{T})\mapsto((\pi_{T})_{*}V_{T},(\pi_{T})_{*}\phi_{T})$ (8.10)
where $\pi_{T}:\Delta_{T}\longrightarrow\mathcal{X}\times_{S}T$ is the modification map and $T$ is an $S$-scheme. It is straightforward to check that the map is $GL_{N(m)}$-equivariant morphism.
8.2 Dimension and local properties of $M^{cl}_{Gie,0}$
In this subsection, we will compute the dimension of $M^{cl}_{Gie,0}$ and show that it is a local complete intersection.
8.2.1 Relative log-symplectic reduction
We will do this by a log-symplectic reduction on every atlas of $N_{Gie,0}$. Let us pick one of the atlas $\mathcal{Y}^{m}_{S}$ of $M^{cl}_{Gie}$. The group $GL_{N(m)}$ acts on $\mathcal{Y}^{m}_{S}$ and the quotient stack $[\mathcal{Y}^{m}_{S}/GL_{N(m)}]$ is an open sub-stack of $N_{Gie}$. We consider the relative log-cotangent bundle $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$. Since the $GL_{N(m)}$ action preserves the normal crossing divisor $\mathcal{Y}^{m}_{0}$ (the closed fibre of $\mathcal{Y}^{m}_{S}\longrightarrow S$), the action of $GL_{N(m)}$ lifts to an action on $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ with a moment map $\mu_{\log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{gl}_{N(m)}^{*}$. Notice that the action of $GL_{N(m)}$ has a generic stabiliser $\mathbb{G}_{m}$; therefore the map $\mu_{\log}$ actually factors through $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$. Since we know that there exists an open subset of $\mathcal{Y}^{m}_{S}$ where the action of $GL_{N(m)}$ has stabiliser isomorphic to $\mathbb{G}_{m}$ (namely, the locus of stable vector bundles), therefore the map $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$ is surjective. Then one can show that $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0)$. To see this, notice that $\mathcal{Y}^{m}_{S}$ is a principal $GL_{N(m)}$-bundle over an open subset of the stack $N_{Gie}$. Let $a$ denote the projection $a:\mathcal{Y}^{m}_{S}\longrightarrow N_{Gie}$. Let us write the cotangent sequence.
$0\longrightarrow a^{*}\Omega^{\log,*}_{N_{Gie}/S}\longrightarrow T^{\log,*}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}$ (8.11)
where the last map is given by $\mu_{\log}$. Notice that $[T^{\log,*}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}]$ is the cotangent complex of the stack $N_{Gie}$. Therefore, using standard arguments from deformation theory of vector bundles it follows that the 00-th cohomology of the complex at a point $[\mathcal{O}^{N(m)}\rightarrow\mathcal{E}]\in\mathcal{Y}^{m}_{S}$ is isomorphic to $H^{1}(\mathcal{E}nd\mathcal{E})^{\vee}\cong\mathsf{Hom}(\mathcal{E},\mathcal{E}\otimes\omega)$. Therefore by Definition 8.7, $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0)$. Therefore $[\mu_{\log}^{-1}(0)/GL_{N(m)}]$ is an open subset of $M^{cl}_{Gie}$. Therefore, we will compute the dimension of $\mu_{\log}^{-1}(0)$ and that it is a local complete intersection. We recall a result on the dimension of the image of a cotangent fibre under the moment map.
Lemma 8.9.
$\dim\mu_{\log}(\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\leavevmode\nobreak\ \dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}$, where $\mathfrak{gl}_{N(m),y}$ is the Lie algebra of the stabilizer of $y\in\mathcal{Y}^{m}_{S}$ under the action of $GL_{N}(m)$.
Proof.
Consider the diagram
[…]
Here $K$ denotes the kernel of the natural map $T^{log}_{\mathcal{Y}^{m}_{S}/S,y}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$. Notice that the map $\mathfrak{gl}_{N(m)}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$ is the differential of the orbit map, and it factors through $T^{log}_{\mathcal{Y}^{m}_{S}/S,y}$ because the action of $GL_{N(m)}$ preserves the normal crossing divisor. It is well known that the rightmost vertical map is injective [37, Lemma 2.4.1]. Therefore we can complete the diagram.
[…]
Therefore we see that $\ker(\iota_{log})\subset\ker(\iota)=\mathfrak{gl}_{N(m),y}$. Since the moment and log-moment maps are the dual of the maps $\iota$ and $\iota_{log}$, therefore $\dim\leavevmode\nobreak\ \mu_{log}(\Omega^{log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\leavevmode\nobreak\ \dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}$. ∎
Proposition 8.10.
The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([37, Definition 2.1.2]) with normal crossing singularities.
Proof.
From [19, Theorem 9.5],it follows that the normalisation of $N^{cl}_{Gie,0}$ is a $KGL_{n}$ bundle over the stack of vector bundles $\mathrm{Bun}(\tilde{X}_{0})$ of rank $n$ and degree $d$ over the normalisation $\tilde{X}_{0}$ of the nodal curve $X_{0}$. Here $KGL_{n}$ denotes the compactification of $GL_{n}$ constructed by Kausz [20]. More precisely, let $E$ is the universal $GL_{n}$ bundle over $\tilde{X_{0}}\times\mathrm{Bun}(\tilde{X_{0}})$. Consider the $GL_{n}\times GL_{n}$ bundle $E_{x_{1}}\times_{\mathrm{Bun}(\tilde{X_{0}})}E_{x_{2}}$ over $\mathrm{Bun}(\tilde{X_{0}})$. Then the associated $KGL_{n}$ fibration $(E_{x_{1}}\times_{Bun(\tilde{X_{0}})}E_{x_{2}})\times_{GL_{n}\times GL_{n}}KGL_{n}\cong\widetilde{N^{cl}_{Gie,0}}$. This is the stack of Gieseker vector bundle data ([19, Definition 4.7]), which is equivalent to ([10, Lemma 5.6]) the stack of marked Gieseker vector bundles (i.e., a Gieseker vector bundle with a marked node ([10, Definition 5.1])). It is obvious that the automorphism group of a marked Gieseker vector bundle is isomorphic to the automorphism group of the corresponding Gieseker vector bundle. Therefore, it is enough to show that the stack of Gieseker vector bundle data is smooth, equidimensional, irreducible and almost very good.
Now let us recall that the map $\tilde{\pi}:\widetilde{N^{cl}_{Gie,0}}\longrightarrow\mathrm{Bun}(\tilde{X}_{0})$ is given by $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\mapsto h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2})$, where $h:X^{m,n}\longrightarrow\tilde{X}_{0}$ is the modification map ([19, Lemma 9.3]). Here $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)$ is a Gieseker vector bundle data and here $\phi$ is not a Higgs bundle but an identification between the fibres ~$\tilde{\mathcal{E}}_{s_{1}}\xrightarrow{\cong}\tilde{\mathcal{E}}_{s_{2}}$ ([19, Definition 4.7]). It is straightforward to check that the induced map
$\operatorname{Aut}(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\longrightarrow\operatorname{Aut}(h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2}))$
is injective. For notational convenience, let us denote $Y:=\widetilde{N^{cl}_{Gie,0}}$ and $Z:=\mathrm{Bun}(\tilde{X}_{0})$.
From [4, Proposition 2.1.2] and [37, Definition 2.1.2, Remark 2.1.3], it follows that $Z$ is smooth, irreducible, almost very good stack. More precisely, the $\operatorname{codim}_{Z}(Z_{k})>k\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\geq 1$, where $Z_{k}:=\{z\in Z\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{dim}\leavevmode\nobreak\ \operatorname{Aut}\leavevmode\nobreak\ (z)=1+k\}$. From the fact that $\operatorname{Aut}(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\subset\operatorname{Aut}(h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2}))$, it follows that $\tilde{\pi}(Y_{k})\subset Z_{k}$. Therefore, $Y_{k}\subset\tilde{\pi}^{-1}(Z_{k})$ and
$\dim\leavevmode\nobreak\ Y_{k}\leq\dim\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\pi}^{-1}(Z_{k})=\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}+\dim\leavevmode\nobreak\ \leavevmode\nobreak\ KGL_{n}.$
Now for all $k>0$ we have
$\begin{split}\operatorname{codim}_{Y}(Y_{k})=&\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y_{k}\\ &\geq\leavevmode\nobreak\ \leavevmode\nobreak\ \dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ KGL_{n}=\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}\\ &=\operatorname{codim}_{Z}(Z_{k})>k.\end{split}$
Therefore, $\widetilde{N^{cl}_{Gie,0}}$ is smooth irreducible, almost very good stack and $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack. ∎
Theorem 8.11.
The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $12\cdot\text{dim}\leavevmode\nobreak\ \leavevmode\nobreak\ N_{Gie,0}+$ 1.
Proof.
It is enough to show that $\mu_{\log,0}^{-1}(0)=\mathcal{Y}^{H,m}_{0}$ is a local complete intersection of dimension $2\cdot\text{dim}\leavevmode\nobreak\ \leavevmode\nobreak\ N_{Gie,0}+1+\dim\leavevmode\nobreak\ GL_{N(m)}$. Here $\mu_{log,0}$ denotes the restriction of $\mu_{\log}$ to the closed fibre of $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow S$. Since the map $\mu_{\log,0}:\Omega^{\log}_{\mathcal{Y}^{m}_{0}}\longrightarrow\mathfrak{pgl}^{*}_{N(m)}$ is surjective, the dimension of the generic fibre is equal to $\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(M)}+1$. Therefore for every irreducible component $I$ of $\mu^{-1}_{\log,0}(0)$, we have
$\dim\leavevmode\nobreak\ I\geq\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(M)}+1.$ (8.14)
Suppose that there exists a component $I\subseteq\mu^{-1}_{\log,0}(0)$ which does not dominates $N_{Gie,0}$. In that case $p(I)\subseteq N_{k}$ for some $k\geq 1$, where $N_{k}:=\{x\in N|\dim\leavevmode\nobreak\ Stab(x)=k+1\}$. A priori, the generic point of $p(I)$ may belong to the singular locus of $N$. But, in any case, using Lemma 8.9, we have
[…]
$<(\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-k)+\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}\hskip 100.00015pt$
$(\text{because}\leavevmode\nobreak\ N_{Gie,0}\leavevmode\nobreak\ \text{ is a almost very good stack})$
$=2\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+(\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m),y}-(k+1))+1*\hskip 50.00008pt$
$=2\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+1\hskip 160.00024pt$
$(\text{because}\leavevmode\nobreak\ \leavevmode\nobreak\ \dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m),y}-(k+1),\text{by definition})$
[…]
where $y$ denotes the generic point of $p(I)$. But this is a contradiction. Therefore, $k$ must be equal to 00, which implies that every component $I$ dominates $\mathcal{Y}^{m}_{0}$ and $\mu^{-1}_{log,0}(0)$ is an equi-dimensional of dimension $12\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+$ 1. Once we know the dimension, it is easy to see that $\mu^{-1}_{log,0}(0)$ is a local complete intersection. Therefore, $M^{cl}_{Gie,0}$ is an equidimensional local complete intersection stack.
We define
$M^{cl,sm}_{Gie,0}:=\{(\mathcal{E},\phi)\in M^{cl}_{Gie,0}\ \ |\ \ \mathbb{H}^{2}(\mathcal{C}(\mathcal{E},\phi))\cong\mathbb{k}\},$
where $\mathcal{C}(\mathcal{E},\phi))$ denote the complex $[\mathcal{E}nd\mathcal{E}\xrightarrow{[-,\phi]}\mathcal{E}nd\mathcal{E}\otimes\omega]$. This substack is obviously an open subset of $M^{cl}_{Gie,0}$ because the minimum dimension of $\mathbb{H}^{2}(\mathcal{C}(\mathcal{E},\phi))$ for any Gieseker-Higgs bundle is $1$. Using the fact that $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack with a generic stabiliser of dimension $1$ and the fact that $M^{cl}_{Gie,0}$ is equi-dimensional, it is not difficult to see that $M^{cl,sm}_{Gie,0}$ is also dense in $M^{cl}_{Gie,0}$.
But notice that $M^{cl,sm}_{Gie,0}$ is a vector bundle over $N_{Gie,0}$ with fibres $H^{0}(\mathcal{E}nd\mathcal{E}\otimes\omega)$ (which remains constant in this locus). Since $N_{Gie,0}$ is irreducible, $M^{cl,sm}_{Gie,0}$ is irreducible. Therefore, $M^{cl}_{Gie,0}$ is also irreducible because $M^{cl,sm}_{Gie,0}$ is dense in $M^{cl}_{Gie,0}$. ∎
Remark 8.12.
Since the geometric fibres of the map $M^{cl}_{Gie}\longrightarrow S$ are of dimension equal to $12n^{2}(g-1)+$ 1, therefore
$\dim M^{cl}_{Gie}=2n^{2}(g-1)+1+\dim S=2n^{2}(g-1)+2.$
Remark 8.13.
As seen in subsubsection 8.2.1, the stack $M^{cl}_{Gie}$ can be covered by open substacks of the form $[\mu_{log}^{-1}(0)/GL_{N(m)}]$, where $\mu_{log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$ is the moment map and $\coprod_{m\geq m_{0}}\mathcal{Y}^{m}_{S}$ is a suitable atlas of the stack $N_{Gie}$. Now notice that the dimension of $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}$ is equal to $2n^{2}(g-1)+2N(m)^{2}+\dim S$. Therefore the expected dimension of $\mu_{\log}^{-1}(0)$ is equal to
$2n^{2}(g-1)+N(m)^{2}+1+\dim S.$
From the previous remark, we know that the dimension of $M^{cl}_{Gie}$ is
$22n^{2}(g-1)+1+\dim S=2n^{2}(g-1)+$ 2
and therefore the dimension of $\mu_{\log}^{-1}(0)$ is equal to $2n^{2}(g-1)+N(m)^{2}+1+\dim S$, which is precisely the expected dimension. Now since $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}$ is a smooth variety, therefore $\mu_{log}^{-1}(0)$ is a local complete intersection. Hence, $M^{cl}_{Gie}$ is also a local complete intersection.